### v1

# On the Compositional Generalization of Multimodal LLMs for Medical Imaging

Abstract
Multimodal large language models (MLLMs) hold significant potential in the medical field, but their capabilities are often limited by insufficient data in certain medical domains, highlighting the need for understanding what kinds of images can be used by MLLMs for generalization. Current research suggests that multi-task training outperforms single-task as different tasks can benefit each other, but they often overlook the internal relationships within these tasks, providing limited guidance on selecting datasets to enhance specific tasks. To analyze this phenomenon, we attempted to employ compositional generalization (CG)—the ability of models to understand novel combinations by recombining learned elements—as a guiding framework. Since medical images can be precisely defined by Modality, Anatomical area, and Task, naturally providing an environment for exploring CG. Therefore, we assembled 106 medical datasets to create Med-MAT for comprehensive experiments. The experiments confirmed that MLLMs can use CG to understand unseen medical images and identified CG as one of the main drivers of the generalization observed in multi-task training. Additionally, further studies demonstrated that CG effectively supports datasets with limited data and delivers consistent performance across different backbones, highlighting its versatility and broad applicability. Med-MAT is publicly available at github.com/FreedomIntelligence/Med-MAT.
UTF8gkai
On the Compositional Generalization of Multimodal LLMs
for Medical Imaging
Zhenyang Cai†, Junying Chen†, Rongsheng Wang†, Weihong Wang, Yonglin Deng, Dingjie Song, Yize Chen, Zixu Zhang, Benyou Wang∗ The Chinese University of Hong Kong, Shenzhen wangbenyou@cuhk.edu.cn
1 Introduction
Multimodal large language models (MLLMs) Liu et al. (2023); Li et al. (2024); Chen et al. (2024b) are showing great promise for the medical community, facilitating efficient consultations for doctors and providing patients with anytime access to their medical conditions. However, limited data on rare or privacy-restricted diseases often restricts MLLMs’ performance, making it crucial to explore what kinds of images can be used by MLLMs to generalize these diseases or learn them efficiently.
Current research on medical image generalization Mo and Liang (2024); Ren et al. (2024) has demonstrated that models trained on multiple tasks outperform those trained on a single task as they can leverage potential knowledge from other tasks. However, these studies have not conducted a detailed analysis of which data within multitask datasets can mutually complement each other. Consequently, we aim to investigate the phenomenon of mutual improvement in MLLMs’ understanding of medical images from the perspective of composition generalization.
Compositional generalization (CG) Xu et al. (2022); Li et al. (2019) refers to a model’s ability to learn fundamental elements and recombine them in novel ways to understand unseen combinations (Figure 1). In medical imaging, each image can be categorized by Modality, Anatomical area, and medical Task, presenting numerous natural opportunities for CG. We defined these three elements as the MAT-Triplet and collected 106 medical datasets, subsequently merging those that share the same MAT-Triplet to create the Med-MAT dataset. Ultimately, Med-MAT comprises 53 subsets, encompassing 11 modalities, 14 anatomical regions, and 13 medical tasks, providing a foundation for investigating CG and other generalization methods.
To verify the existence of CG, we designated certain datasets as Target data and selected all the Related data from Med-MAT that shared the same MAT-Triplet with the Target data for generalization. In addition, we conducted larger-scale experiments and observed the changes in model generalization performance after deliberately disrupting CG. Ultimately, the experiments demonstrated that MLLMs can leverage CG to understand unseen medical images and CG is one of the key factors driving the generalization observed in multi-task training.
Subsequently, the practical applications of CG and its effects on different backbones were further explored. We found that increasing the volume of CG combinations continuously enhanced the model’s understanding of the Target data. While CG’s direct impact on unseen data was not always apparent, CG could assist the model in efficient fitting with limited data by being incorporated into the training data of the Target. Furthermore, we discovered that CG exists in different MLLM backbones, underscoring CG’s broad applicability.
Here are the key contributions of our work: 1) A VQA dataset, Med-MAT, has been constructed, providing a platform to explore the generalization of MLLMs on medical images. 2) Through this dataset, we observed that MLLMs can utilize compositional generalization to understand unseen images, demonstrating that this is one of the main forms of generalization for medical MLLMs. 3) Finally, the contribution of CG to understanding medical domains with limited data has been further explored, with its presence across different backbones demonstrating its broad applicability.
| Model | 02 | 03 | 07 | 08 | 09 | 11 | 13 | 14 | 15 | 16 | 18 | 19 | 21 | 22 | 23 | 25 | 26 | 28 | 30 | 31 | 32 | 33 | 35 | 36 | 37 |
| Baseline | 22 | 47 | 40 | 25 | 26 | 27 | 28 | 24 | 22 | 24 | 25 | 23 | 49 | 26 | 25 | 24 | 49 | 30 | 49 | 21 | 49 | 20 | 25 | 23 | 19 |
| Single-task Training | 24 | 49 | 50 | 68 | 65 | 76 | 83 | 53 | 61 | 32 | 29 | 26 | 57 | 53 | 28 | 24 | 57 | 64 | 89 | 60 | 97 | 54 | 29 | 51 | 49 |
| Multi-task Training | 96 | 89 | 80 | 80 | 79 | 97 | 92 | 88 | 76 | 57 | 88 | 74 | 87 | 86 | 93 | 52 | 98 | 72 | 94 | 61 | 100 | 72 | 75 | 60 | 50 |
| Model | 01 | 04 | 05 | 06 | 10 | 12 | 17 | 20 | 24 | 27 | 29 | 34 |
| Baseline | 32 | 25 | 33 | 33 | 48 | 27 | 33 | 13 | 34 | 37 | 31 | 20 |
| Multi-task Training | 39 | 26 | 70 | 31 | 58 | 38 | 61 | 40 | 35 | 41 | 55 | 50 |
2 Med-MAT
Most existing datasets for MLLMs Zhang et al. (2023c); Li et al. (2024); Chen et al. (2024b), primarily VQA datasets, provide broad coverage but lack attribute annotations for individual samples, which are not suitable for CG exploration. To address this gap, we curated a large collection of image-text pairs to develop Med-MAT, ensuring that each sample is explicitly defined by MAT-Triplet. Section 2.1 provides data processing methods; and Section 2.2 presents a pilot experiment.
2.1 Data Processing
Data Construction: Med-MAT contains a total of 106 image-label pair medical datasets, sourced from various medical public challenges or high-quality annotated datasets. All datasets are categorized according to their MAT-Triplet, with data having identical elements grouped into a single subset (Figure 2). Labels are manually clustered to ensure that annotations with the same meaning are not repeatedly used. In total, Med-MAT covers 11 medical modalities, 14 anatomical areas, and 13 medical tasks, hoping that it can spread across various medical tasks like a mat. (Data lists are shown in Appendix B)
Data Distribution: All subsets are divided into training and test sets following their original distributions or using a 9:1 ratio. To ensure a fair comparison, each training set is limited to 3,000 samples, with label balance maintained as much as possible. Any subset that cannot meet this requirement is treated as an OOD (out-of-distribution) dataset. For the test sets, we strictly balance the number of samples per label to ensure that the accuracy metric reliably reflects model performance.
QA Pairs Construction: To enable MLLMs to directly train and test on Med-MAT, all image-label paired data were converted into a visual question-answering (VQA) format (Figure 3). Specifically, each subset was manually assigned 6 instructions to guide the MLLM in answering the subset task. For convenience, all samples were converted into single-choice questions with up to four options, and the remaining distractor options were randomly drawn from other labels within the subset. To mitigate potential evaluation biases arising from varying option counts, the ImageWikiQA dataset Zhang et al. (2024b), a non-medical dataset consisting of single-answer, four-option questions, was incorporated during the training.
2.2 A Pilot Study of Data Composition
Experiments Setup: Experiments in this paper mainly focused on classification datasets to explore the model’s combinatorial generalization capabilities across various image types. The base model LLaVA-v1.5-7B-Vicuna Liu et al. (2023) features a transparent pretraining process and uses very little medical data during training, thereby minimizing the risk of knowledge leakage. By leveraging the flexibility of MLLM, we achieved task switching and generalization simply by modifying the corresponding prompts, significantly simplifying generalization studies. Each experiment was conducted over 5 epochs using 8 A800 (80GB) GPUs, with a batch size of 32 and a learning rate of 5e-6. Generalization performance was assessed by comparing the model’s accuracy on the Target data.
Pilot Study: To explore the benefits of data composition for downstream tasks, all in-distribution (ID) datasets were combined for multi-task training, with single-task training on individual ID datasets serving as the control. The results (Table 1 and 2) showed that multi-task training outperformed single-task training on specific tasks, and also surpassed the baseline in predicting unseen OOD datasets. This suggests that certain data combinations can enhance the model’s classification performance, highlighting the potential for CG. Exploring which combinations provide valuable insights for downstream medical tasks in a promising avenue for further research.
Take-away 1: Certain data combinations can assist MLLMs in medical classification tasks.
3 Proof of Concept on CG
To explore whether MLLMs can leverage CG to understand unseen images, controlled variable studies were first conducted using all possible CG combinations to determine if CG exists among the MAT-Triplet (Section 3.1). Subsequently, the number of combinations was increased to explore more generalizable insights, aiming to assess the performance gains of multi-task training (Section 3.2).
| Related Combination | Target Subset | Baseline | Baseline+ | Trained | |
| Lung, COVID | Brain, Cancer | Lung, Cancer | 25 | 25 | 27 |
| Lung, Cancer | Brain, State | Lung, State | 47 | 46 | 50 |
| Brain, Cancer | Lung, State | Brain, State | 33 | 50 | 57 |
| Bones, Level | Lung, State | Bones, State | 49 | 53 | 51 |
| Bones, Level | Brain, State | Bones, State | 49 | 53 | 72 |
| Bones, Level | Breast, Diseases | Bones, Diseases | 37 | 33 | 39 |
| Bones, Level | Lung, Diseases | Bones, Diseases | 37 | 33 | 43 |
| Bones, Level | Chest, Diseases | Bones, Diseases | 37 | 31 | 43 |
| Bones, State | Breast, Diseases | Bones, Diseases | 37 | 37 | 43 |
| Bones, State | Lung, Diseases | Bones, Diseases | 37 | 37 | 43 |
| Bones, State | Chest, Diseases | Bones, Diseases | 37 | 37 | 41 |
| Lung, COVID | Breast, Diseases | Lung, Diseases | 49 | 48 | 51 |
| Lung, COVID | Bones, Diseases | Lung, Diseases | 49 | 48 | 52 |
| Lung, COVID | Chest, Diseases | Lung, Diseases | 49 | 48 | 51 |
| CT, Cancer | X-ray, COVID | CT, COVID | 47 | 46 | 72 |
| CT, COVID | X-ray, Diseases | X-ray, COVID | 30 | 21 | 49 |
| CT, State | X-ray, Diseases | X-ray, State | 30 | 21 | 46 |
| CT, State | X-ray, Cancer | CT, Cancer | 33 | 28 | 28 |
| CT, Brain(State) | X-ray, Bones | X-ray, Brain | 49 | 49 | 91 |
| CT, Brain | X-ray, Lung | X-ray, Brain | 49 | 50 | 81 |
| CT, Brain(Cancer) | X-ray, Bones | X-ray, Brain | 25 | 51 | 74 |
| CT, Brain | X-ray, Lung | X-ray, Brain | 49 | 52 | 52 |
| X-ray, Brain | CT, Lung(State) | CT, Brain(State) | 33 | 50 | 60 |
| X-ray, Lung | CT, Brain | CT, Lung(Cancer) | 25 | 25 | 36 |
| X-ray, Lung | CT, Brain(State) | CT, Lung | 47 | 50 | 81 |
| X-ray, Lung | CT, Brain(Cancer) | CT, Lung | 47 | 50 | 71 |
| CT, Lung (State) | X-ray, Bones | X-ray, Lung | 30 | 32 | 28 |
| CT, Lung (State) | X-ray, Brain | X-ray, Lung | 30 | 32 | 35 |
| CT, Lung (Cancer) | X-ray, Bones | X-ray, Lung | 30 | 32 | 41 |
| CT, Lung (Cancer) | X-ray, Brain | X-ray, Lung | 30 | 32 | 42 |
| Der, Skin, Cancer | FP, Fundus, Diseases | Der, Skin, Diseases | 25 | 29 | 33 |
| Der, Skin, Cancer | OCT, Retine, Diseases | Der, Skin, Diseases | 25 | 29 | 33 |
| Der, Skin, Diseases | DP, Mouth, Cancer | Der, Skin, Cancer | 40 | 33 | 63 |
| Der, Skin, Diseases | Mic, Cell, Cancer | Der, Skin, Cancer | 40 | 33 | 63 |
| DP, Mouth, State | Der, Skin, Cancer | DP, Mouth, Cancer | 48 | 50 | 52 |
| DP, Mouth, State | Mic, Cell, Cancer | DP, Mouth, Cancer | 48 | 50 | 55 |
| FP, Fundus, Diseases | Mic, Cell, Level | FP, Fundus, Level | 33 | 36 | 42 |
| Mic, Cell, Cell Identification | FP, Fundus, Level | Mic, Cell, Level | 23 | 33 | 32 |
| Mic, Cell, Cell identification | Der, Skin, Cancer | Mic, Cell, Cancer | 49 | 50 | 50 |
| Mic, Cell, Cell identification | DP, Mouth, Cancer | Mic, Cell, Cancer | 49 | 51 | 62 |
| Mic, Cell, Level | Der, Skin, Cancer | Mic, Cell, Cancer | 49 | 51 | 52 |
| Mic, Cell, Level | DP, Mouth, Cancer | Mic, Cell, Cancer | 49 | 51 | 58 |
| Mic, Cell, Cancer | FP, Fundus, Level | Mic, Cell, Level | 23 | 24 | 27 |
3.1 Controlled Variable Study on CG
3.1.1 Experiment Setup
In this section, the existence of CG was explored from a finer perspective, focusing on CG with only two elements varying while the third remained constant. For example, we fixed the Modality and examined the CG between Areas and Tasks. Additionally, we identified specific Modality-Area pairs, such as dermoscopy paired consistently with skin, which were treated as a special category. These 4 different fixed formats were classified into distinct Direction Types.
Following the setup in Section 2.2, we evaluated the model’s performance on Target data. Baseline refers to the model without any training, while Trained refers to the model trained solely on Related data. To ensure that our conclusions are not influenced by the amount of training data, we randomly sampled an equal number of data from the Unrelated subsets, and this configuration is referred to as Baseline+.
3.1.2 Analysis of CG in MLLMs
Results are shown in Table 3 and it can be observed that almost all CG combinations are able to generalize to downstream tasks, highlighting that MLLMs can leverage CG to generalize Target data across all Direction Types. Besides that, since this experiment focused solely on two-element tuples, we further investigated three-element tuples in Section 4.3, where we also observed similarly strong generalizations when obtaining MAT-Triplet elements from three different datasets.
Take-away 2: MLLMs can leverage CG to understand unseen medical images.
In the Baseline+ setting, we removed all datasets sharing any MAT-Triplet element with the Target data. Consequently, Baseline+ models perform at near-random levels on the test set, indicating they failed to acquire target-relevant knowledge. This suggests that only datasets related through the MAT-Triplet can help the model learn and generalize to new target tasks.
Take-away 3: Generalization only arises from MAT-Triplet related medical datasets.
3.1.3 Analysis of Non-obvious Cases
Some Trained models show minimal gains or performance declines, especially in Level or Diseases Identification tasks. The images in the Level Identification are very similar, while the Diseases Identification involves a large number of target labels, both of which require large datasets for training, which may explain the poor performance. To further analyze them, we modified the evaluation scenario to use CG combinations for assisting in Target data training with limited samples (Section 4.2).
3.2 Scaling the Combination Number of CG
To ensure the universality of CG, we expanded the number of combination datasets to evaluate its effects on a larger scale. This section explores the following questions: (Q1) While Meta CG experiments indicate that Unrelated combinations provide no benefit to Target data, can generalization arise when training incorporates more Unrelated combinations, simulating a multi-task scenario? (Q2) Previous studies suggest that multi-task training generally promotes better generalization than single-task training. To what extent does this generalization originate from CG?
3.2.1 Experiment Setup
Selection Strategy: To ensure a balanced evaluation of Related and Unrelated combinations, Subset 03 and Subset 28 were chosen as Target datasets because they exhibit the most balanced ratios of Related to Unrelated subsets (13:11 for Subset 03 and 11:13 for Subset 28), making them ideal for providing a diverse range of compositions in the scale-up experiments.
The baseline was trained on all subsets excluding the Target data to evaluate the claim that mixing multi-task data enhances generalization (All Data). To construct multiple comparative experiments, models were further trained on either Related or Unrelated subsets (All Related / All Unrelated) to address Q1. For Q2, individual MAT-Triplet elements were systematically removed from the Related subsets (Related w/o Modality / Area / Task), disrupting CG and assessing the ability to maintain generalization. To ensure consistency, the total data volume in all experiments was limited to 15,000 samples, aligning with the number of ID subsets available after excluding related tasks from Subset 03.
3.2.2 Analysis of Scaling Experiment
Figure 4 illustrates the results. It can be observed that even when we expanded the Unrelated combination volumes and increased task diversity, the performance of All Unrelated remains close to the Baseline, indicating that these datasets can not support MLLMs to understand the Target data.
Take-away 4: Datasets without MAT-Triplet overlap offer limited benefit for generalization even in the multi-task training scenario (Q1).
Besides, w/o Modality / Area / Task showed significant accuracy drops compared to All Related, despite holding the training data volume constant. Notably, All Related performs at a level comparable to All Data, where all datasets are included in training. This suggests that CG provides significant help for the generalization brought by multi-task training.
Take-away 5: CG is one of the primary forms of generalization for MLLMs trained on multi-task dataset in medical imaging (Q2).
| Related Combination | Target Subset | Baseline | Trained | ||
| CT - Subset02 | Brain - Subset22 | Cancer - Subset07 | CT, Brain, Cancer | 28 | 26 |
| CT - Subset03 | Brain - Subset22 | Cancer - Subset21 | CT, Brain, Cancer | 28 | 25 |
| CT - Subset02 | Brain - Subset22 | State - Subset09 | CT, Brain, State | 33 | 64 |
| CT - Subset03 | Brain - Subset22 | State - Subset26 | CT, Brain, State | 33 | 70 |
| X-ray - Subset25 | Lung - Subset03 | Diseases - Subset02 | X-ray, Lung, Diseases | 30 | 45 |
| X-ray - Subset26 | Lung - Subset03 | Diseases - Subset02 | X-ray, Lung, Diseases | 30 | 38 |
| X-ray - Subset26 | Lung - Subset03 | Diseases - Subset08 | X-ray, Lung, Diseases | 30 | 44 |
| X-ray - Subset26 | Breast - Subset24 | Diseases - Subset02 | X-ray, Breast, Diseases | 31 | 32 |
| X-ray - Subset28 | Breast - Subset24 | Diseases - Subset08 | X-ray, Breast, Diseases | 31 | 52 |
4 In-depth Analysis of CG
To further explore the application and applicability of CG, several research questions were proposed in this section. In order to present rich results with minimal consumption of computational resources, specific data Selection Strategies were applied for each question. Here are the research questions:
(RQ 1): How does the quantity of Related data affect generalization, and should it be maximized to ensure CG quality?
(RQ 2): Can CG combinations help the model efficiently fit Target data when only a small amount of Target data is available?
(RQ 3): Previous studies demonstrated that CG exists between two MAT-Triplet elements, does CG still exist if the three MAT-Triplet elements come from three different datasets?
(RQ 4): Does CG exist between images from Detection tasks and Classification tasks?
(RQ 5): Can different MLLM backbones leverage CG to interpret unseen medical images?
4.1 The Relationship Between CG Performance and Data Volume
Selection Strategy: This section focuses on examining how the data volume of Related combinations influences generalization. To highlight the generalization trends, the combinations with strong generalization results were selected from the main experiments. For fairness, we chose the combinations across four types where Trained results exceed both Baseline and Baseline+ by at least 10. If multiple combinations meet the criteria, a random seed of 42 was used to determine the selection.
In this experiment, the amount of combination data was gradually increased (from 0 to 750, 1500, 2250, and 3000), and the trend of the model’s accuracy on Target data after training was observed. The experimental results are shown in Figure 5, where the red line represents the accuracy curve gained from Related combinations, and the purple line represents the gain from Unrelated combinations. Models trained on Related combinations demonstrate strong performance across all four Target data sets. As the data volume of Related combinations increases, the model’s understanding of Target data improves consistently.
Take-away 6: Generalization effect improves with increasing CG combination’s volume. (RQ 1)
4.2 Achieving Data-Efficient Training through CG
Selection Strategy: This experiment evaluates how CG aids training and dataset fitting with limited Target data. To highlight performance trends, four combinations with poor generalization (marked in red in Table 3) were selected across the four Direction Types, ensuring both models in the control group started with similar baseline accuracies. When multiple poor-performing combinations existed within a Direction Type, the more challenging datasets were chosen based on Table 1 to better observe potential improvements.
In this experiment, the amount of Target data was progressively increased (from 0 to 500, 1000, 1500, and 2000), and a fixed number of combination data were incorporated during training. Figure 6 displays the results, with red lines showing accuracy gains from Related combination and purple lines from Unrelated ones. It can be observed that training with Related combination data reaches peak performance more quickly, demonstrating another form of CG. Therefore, although CG performs poorly on these data due to task characteristics, it still helps the model quickly adapt with only a small amount of Target data.
Take-away 7: Although CG does not always provide direct generalization gains, it can help Target data achieve data-efficient training. (RQ 2)
4.3 CG with All MAT-Triplet Elements from Different Sources
In previous controlled experiments (Section 3.1), one element of the MAT-Triplet was kept constant while CG was explored in the remaining two elements. To ensure that all the 3 MAT-Triplet elements of the target data originated from three distinct datasets, additional experiments were conducted to further validate the effectiveness of CG. For these experiments, all possible combinations meeting the criteria in Med-MAT were selected (Selection Strategy). The results presented in Table 4 demonstrate that most combinations can effectively generalize to the Target data.
Take-away 8: CG remains effective when all elements come from different datasets (RQ 3).
4.4 Detection-Capable MLLMs can Leverage Detection Data for CG
Previous studies Ren et al. (2024); Wang et al. (2025) have demonstrated that jointly training classification and detection tasks can mutually enhance their performance. Building on this, we explored CG within this context. To ensure comprehensive conclusions, we conducted experiments using Next-Chat and MiniGPT-v2, which represent the two mainstream approaches for MLLMs to perform segmentation tasks. The former treats bounding boxes as embeddings and decodes them into coordinates using an additional visual decoder, while the latter processes coordinate points as special text tokens and generates bounding box coordinates directly as output text. The final results (Figure 7) show that all CG combinations demonstrated the model’s successful utilization of detection data for CG to the Target data.
Take-away 9: MLLMs can combine knowledge from detection tasks with classification tasks to improve classification accuracy (RQ 4).
4.5 CG Exists Across Different MLLM Backbones
LLaVA was selected as the baseline because its training data and processes are publicly available, ensuring minimal exposure to medical images and preventing bias in the integration of medical image knowledge into the MLLM. To verify that CG is not limited to a single framework, we also selected subsets for experiments on Qwen2-VL-7B Wang et al. (2024a) and Llama3.2-11B-Vision Meta AI (2024). The results in Figure 7 demonstrate CG effects similar to those observed in the LLaVA.
Take-away 10: CG persists across different MLLM backbones (RQ 5).
5 Related Work
Medical MLLMs: Recently, adapting MLLMs to medical tasks has gained prominence due to their success in capturing complex visual features. Current MLLMs typically pair a visual encoder with a text-only LLM, aligning image data with language understanding. Such as Med-Flamingo (Moor et al., 2023) and Med-PaLM (Tu et al., 2024), fine-tuned general multimodal models and achieved notable results. Med-Flamingo enhanced OpenFlamingo-9B (Chen et al., 2024a) with medical data, while Med-PaLM adapted PaLM-E (Driess et al., 2023) using 1 million data points. Similarly, LLaVA-Med Li et al. (2024), Med-Gemini (Saab et al., 2024), and HuatuoGPT-Vision Chen et al. (2024b) utilized specialized datasets and instruction tuning to refine medical VQA tasks.
Generalization on Medical Imaging: Generalization in medical imaging Matta et al. (2024) has been extensively studied. Early methods utilized data manipulation techniques, such as data augmentation Li et al. (2022); Zhang et al. (2022), to enhance model generalization on unseen medical data by adapting to varying distributions. Later approaches focused on representation learning Le-Khac et al. (2020), preserving essential image information to enable models to handle more complex scenarios. Additionally, some studies Ren et al. (2024) explore multiple aspects of medical image processing, examining how classification and segmentation tasks can mutually benefit each other.
Detection with MLLMs: MiniGPT-4 Zhu et al. (2023) and LLaVA Liu et al. (2023) enhance LLMs by fine-tuning models on synthetic multimodal instruction data. However, these models are limited to image and text inputs and cannot generate detection bounding boxes or segmentation masks. Recent works have addressed these limitations using various strategies, such as encoding regions as features to allow models to accept regions as input Zhang et al. (2023b), representing object bounding box coordinates with text tokens Wang et al. (2024c); Peng et al. (2023); Chen et al. (2023b), and employing unique identifiers for task instructions to improve learning efficiency. Additionally, some approaches introduce special tokens to represent images and use their hidden states to decode position information Zhang et al. (2023a, 2024a).
6 Conclusion
To explore whether MLLMs can leverage CG to generalize unseen medical data, we constructed the Med-MAT dataset as the research platform for generalization experiments. The results confirmed the existence of CG and identified it as one of the primary factors behind the generalization capability of MLLMs in multi-task learning. Subsequent experiments demonstrated that increasing the volume of CG combinations consistently improved their effectiveness. Furthermore, CG enables medical tasks to efficiently fit with limited data, reducing dependence on the volume of training data. Importantly, CG is present across various MLLM backbones and can even facilitate generalization using data from tasks such as detection, underscoring its broad applicability.
Limitations
The experiment confirms that MLLMs can use CG to understand unseen medical images and achieve data-efficient training. However, as shown in the results in 3.2, even after disrupting CG, the model’s generalization performance declines but still maintains some level of effectiveness. Thus, CG represents only one form of generalization for MLLMs in medical imaging. To support the exploration of other generalization mechanisms, we will make datasets publicly available according to the license.
This study is focused exclusively on medical scenarios, but we believe that similar generalization effects may also apply in other multimodal tasks. Additionally, more granular medical generalization strategies, such as leveraging existing pneumonia data to generalize the detection of newly emerging types of pneumonia, are also promising areas for further investigation.
Potential Risks
Our research focuses on the compositional generalization of MLLMs on medical images, using data sourced from medical challenges and open-source datasets. However, further experiments are needed to mitigate potential risks when deploying this concept in real-world medical settings.
Appendix A More Experiments
A.1 CG on Detection Tasks (RQ 4)
| Related Combination | Target Subset | Baseline | Trained | |
| Lung, Lung Det | Bones, Diseases | Lung, Diseases | 49 | 52 |
| Lung, Lung Det | Breast, Diseases | Lung, Diseases | 49 | 54 |
| Bones, Spinal Error Det | Breast, Diseases | Bones, Diseases | 20 | 30 |
| Bones, Spinal Error Det | Lung, Diseases | Bones, Diseases | 20 | 33 |
| MRI, Diseases Det | End, Level | End, Diseases | 24 | 27 |
| X-ray, Lung Det | CT, COVID | X-ray, COVID | 23 | 26 |
| Der, Skin, Cancer Det | FP, Fundus, Diseases | Der, Skin, Diseases | 24 | 29 |
| Mic, Cell, Cancer Det | CT, Kidney, Diseases | Mic, Cell, Diseases | 24 | 26 |
| Related Combination | Target Subset | Baseline | Trained | |
| Lung, Lung Det | Bones, Diseases | Lung, Diseases | 41 | 47 |
| Lung, Lung Det | Breast, Diseases | Lung, Diseases | 41 | 49 |
| Bones, Spinal Error Det | Breast, Diseases | Bones, Diseases | 31 | 35 |
| Bones, Spinal Error Det | Lung, Diseases | Bones, Diseases | 31 | 37 |
| MRI, Diseases Det | End, Level | End, Diseases | 24 | 26 |
| X-ray, Lung Det | CT, COVID | X-ray, COVID | 22 | 23 |
| Der, Skin, Cancer Det | FP, Fundus, Diseases | Der, Skin, Diseases | 27 | 30 |
| Mic, Cell, Cancer Det | CT, Kidney, Diseases | Mic, Cell, Diseases | 20 | 24 |
Previous studies have shown that training on detection datasets can improve model performance in classification tasks. Building on this insight, we further examine its applicability to CG. For this investigation, all possible detection subset combinations in Med-MAT were selected to explore their impact on CG (Selection Strategy).
Experimental Setup: We conducted generalization experiments for detection and classification. Specifically, we performed generalization validation on Next-Chat Zhang et al. (2023a) and MiniGPT-v2 Chen et al. (2023a). Next-Chat models the bounding box as an embedding and utilizes a decoder for decoding, while MiniGPT-v2 treats the bounding box as a text token, which are common approaches used by existing MLLM implementations for detection. By conducting CG validation using distinct bounding box modeling methods, we further demonstrate the broad applicability of the CG approach. Each experiment was conducted on 8 A800 (80GB) GPUs.
The two backbone was trained separately in this experiment. For Next-Chat, we directly trained the model in its second training stage and fine-tuned it for 2 epochs on our composition datasets with a learning rate of 2e-5, keeping all other training parameters at their default settings. Similarly, for MiniGPT-v2, we trained the backbone model from the second stage, starting with a learning rate of 2e-5 and gradually reducing it to 2e-6 over 3 epochs.
The experiments in Table 5 and Table 6 present the results for Next-Chat and MiniGPT-v2, respectively. The separation regions closely align with the 4 Direction Types defined in the classification task, except for Fix Task, which is excluded due to the substantial differences between detection and classification tasks, making it impossible to find the fix one. The results demonstrate that joint training on detection and classification tasks effectively compositional generalizes to classification tasks in both backbones, leading to higher model accuracy. This indicates that CG is also present in these tasks.
| Related Combination | Target Subset | Baseline | Trained | |
| Bones, State | Breast, Diseases | Bones, Diseases | 61 | 65 |
| Lung, COVID | Bones, Diseases | Lung, Diseases | 80 | 91 |
| CT, COVID | X-ray, Diseases | X-ray, COVID | 35 | 40 |
| CT, State | X-ray, Diseases | X-ray, State | 35 | 43 |
| X-ray, Lung | CT, Brain(Cancer) | CT, Lung | 32 | 33 |
| X-ray, Lung | CT, Brain | CT, Lung(Cancer) | 65 | 72 |
| FP, Fundus, Diseases | Mic, Cell, Level | FP, Fundus, Level | 48 | 45 |
| Mic, Cell, Cell Identification | FP, Fundus, Level | Mic, Cell, Level | 34 | 41 |
| Related Combination | Target Subset | Baseline | Trained | |
| Bones, State | Breast, Diseases | Bones, Diseases | 52 | 59 |
| Lung, COVID | Bones, Diseases | Lung, Diseases | 64 | 75 |
| CT, COVID | X-ray, Diseases | X-ray, COVID | 33 | 38 |
| CT, State | X-ray, Diseases | X-ray, State | 33 | 41 |
| X-ray, Lung | CT, Brain(Cancer) | CT, Lung | 31 | 29 |
| X-ray, Lung | CT, Brain | CT, Lung(Cancer) | 49 | 57 |
| FP, Fundus, Diseases | Mic, Cell, Level | FP, Fundus, Level | 55 | 61 |
| Mic, Cell, Cell Identification | FP, Fundus, Level | Mic, Cell, Level | 10 | 32 |
A.2 Other MLLMs (RQ 5)
The choice to use models like LLaVA, MiniGPT-v2 and NExT-GPT is because their training data and overall process are publicly available, ensuring that they have not been too much exposed to medical images, thus preventing bias in the medical image knowledge injection into the MLLM. To ensure the experiment results are not influenced by the model choice, we also tested several other models on some subsets of Med-MAT and observed similar results.
Selection Strategy: For testing, some generalized combinations were selected from classification tasks 3. Using a random seed of 42, we shuffled each Direction Type’s combinations and selected the first two compositions as test data.
Experimental Setup: We conducted experiments to evaluate the compatibility of CG across different backbone architectures. We selected two MLLMs with representative architectures, namely Qwen2-VL-7B-Instruct Wang et al. (2024b) and Llama-3.2-11B-Vision-Instruct Meta AI (2024), to assess the performance of CG on these models. Each experiment involved full-parameter fine-tuning of all models over 5 epochs, utilizing 8 A800 (80GB) GPUs. The training was performed with a batch size of 32 and a learning rate set to 2e-6, ensuring that all parameters were updated to optimize the model performance.
The experimental results presented in Table 7 and Table 8 demonstrate that the CG persists across these different backbone architectures. This observation indicates that CG is not confined to a specific model type, thereby highlighting its universal applicability and robustness across various model frameworks. Such findings underscore the versatility of CG, suggesting it can be effectively integrated into a wide range of models.
Appendix B The Dataset: Med-MAT
This section provides an overview of Med-MAT. First, a detailed explanation of MAT-Triplet will be presented in B.1. Next, the methods for constructing the QA formatting will be discussed in B.2. Finally, the data composition details and open-source specification will be provided in B.3.
B.1 Details of MAT-Triplet
MAT-Triplet stands for Medical Modality, Anatomical Area, and Medical Task. We define all samples in Med-MAT using these three components and integrate datasets with identical triplets into subsets.
Medical Modality refers to different types of techniques or methods used in medical imaging or data acquisition. Each modality is designed to present the human body’s structures or pathological features in unique ways, providing auxiliary support for clinical diagnosis and treatment. Most modalities exhibit significant visual differences, making them easily distinguishable. Med-MAT encompasses 11 modalities, including common ones such as Computed Tomography (CT), Magnetic Resonance Imaging (MRI), X-ray, Fundus Photography (FP), Endoscopy (End), Optical Coherence Tomography (OCT), and Ultrasound (US), as well as rare and specialized modalities like Colonoscopy (Co), Dermoscopy (Der), Digital Pathology (DP), and Microscopy (Mic).
Anatomical Area refers to specific anatomical structures or regions within the human body or other organisms, defined by distinct anatomical characteristics to describe various body parts, their functions, and relative positions. Med-MAT encompasses 14 anatomical areas, including the cervix, kidney, lung, brain, intestine, bladder, fundus, retina, breast, bones, and chest. To further facilitate data description, additional categories such as skin, mouth, and cell are included as specialized anatomical areas.
Medical Task refers to the specific detection task that needs to be performed on the dataset. Med-MAT includes 13 distinct tasks, with classification tasks encompassing Quality Identification (image quality analysis), COVID Diagnosis, Cancer Diagnosis (determining the presence of a specific disease), State (such as identifying brain hemorrhage), Level Identification (assessing disease severity), and Multiple Classification (classifying multiple diseases or cell types). Given the limited options of COVID Diagnosis and Cancer Diagnosis, these tasks can be interpreted as identifying whether a patient is in a diseased state. To enhance generalization and provide more diverse examples, these tasks are grouped under the broader category of State. In addition, we have 16 datasets defining segmentation or classification tasks with different objectives.
B.2 QA construction method
A large amount of image-label datasets was collected to build the Med-MAT dataset. To ensure compatibility with MLLM training inputs and outputs, all data is transformed into a question-answering format. Questions are formulated based on modality, anatomical area, and medical task, with 6 question prompts applied to each subset.
The labels within each data subset will be clustered to prevent redundant definitions of the same condition. Then, all training set and test set will be converted into multiple-choice questions following the template in Table 8. Each question will have up to four options, with distractor options randomly selected from the corresponding subset.
B.3 Data composition and Open-source Specification
Med-MAT is composed of multiple datasets. After being transformed into different QA formats, the new data is organized into several subsets to support generalization experiments in medical imaging. Table 9 shows all of our subset datasets, which are separated based on different combinations in MAT-Triplet. The specific MAT-Triplets are listed, along with the labels corresponding to the image-label datasets for each subset. Correspondingly, all the image-label datasets are also displayed in Table 10, which includes their names, descriptions of the tasks performed, download links, and the level of accessibility.
All question-answering text datasets in Med-MAT will be publicly available. To accommodate varying access permissions, we will release datasets based on their respective licenses: openly accessible datasets will be directly available, while restricted datasets can be accessed by applying through the links provided in this paper. We hope this dataset will support and advance future generalization experiments on medical imaging.
| Subset No. | Modality | Anatomical Area | Task | Datasets No. |
| 01 | Co | Cervix | Cervical Picture Quality Evaluation | 1 |
| 02 | CT | Kidney | Kidney Diseases Classification | 2 |
| 03 | CT | Lung | COVID-19 Classification | 3,4,6 |
| 04 | CT | Lung | Lung Cancer Classification | 5 |
| 05 | CT | Brain | Brain Hemorrhage Classification | 7 |
| 06 | CT | Brain | Brain Cancer Classification | 8 |
| 07 | Der | Skin | Melanoma Type Classification | 10 |
| 08 | Der | Skin | Skin Diseases Classification | 9, 11-15, 71, 72, 74 |
| 09 | DP | Mouth | Teeth Condition Classification | 16 |
| 10 | DP | Mouth | Oral Cancer Classification | 17 |
| 11 | End | Intestine | Intestine Cleanliness Level | 18 |
| 12 | End | Bladder | Cancer Degree Classification | 19 |
| 13 | End | Intestine | Intestine Diseases Classification | 20 |
| 14 | FP | Fundus | Eye Diseases Classification | 21-23, 26-28, 31, 32, 75 |
| 15 | FP | Fundus | Multiple-labels Eye Diseases Classification | 24, 25, 68 |
| 16 | FP | Fundus | Blindness Level | 29 |
| 17 | FP | Fundus | Retinal Images Quality Evaluation | 30 |
| 18 | Mic | Cell | Cell Type Classification | 33, 36-38, 39-41, 44, 65, 70 |
| 19 | Mic | Cell | Prostate Cancer Degree Classification | 34 |
| 20 | Mic | Cell | Multiple-labels Blood Cell Classification | 35 |
| 21 | Mic | Cell | Cancer Classification | 42, 67 |
| 22 | MRI | Brain | Head Diseases Classification | 44, 45 |
| 23 | OCT | Retina | Retina Diseases Classification | 46, 47 |
| 24 | US | Breast | Breast Cancer Classification | 48 |
| 25 | X-ray | Bones | Degree Classification of Knee | 49, 53 |
| 26 | X-ray | Bones | Fractured Classification | 50, 51 |
| 27 | X-ray | Bones | Vertebrae Diseases Classification | 52 |
| 28 | X-ray | Lung | COVID-19 and Pneumonia Classification | 54-57, 60, 62, 81 |
| 29 | X-ray | Breast | Breast Diseases Classification | 58, 78 |
| 30 | X-ray | Lung | Tuberculosis Classification | 59, 79 |
| 31 | X-ray | Chest | Multiple-labels Chest Classification | 61, 73, 76, 77, 80, 85, 87 |
| 32 | X-ray | Brain | Tumor Classification | 63 |
| 33 | Mic | Cell | Multi-labels Diseases | 84 |
| 34 | FP | Fundus | Level Identification | 66 |
| 35 | X-ray | Bones | Level Identification | 69 |
| 36 | X-ray | Bones | Spinal lesion Classification | 86 |
| 37 | X-ray | Breast | Multi-labels Diseases | 82 |
| 38 | Der | Skin | Lesion Det/Seg | 88-91 |
| 39 | End | Intestine | PolyP Det/Seg | 92-93 |
| 40 | End | Intestine | Surgical Procedures Det/Seg | 94 |
| 41 | End | Intestine | Multi-labels Det/Seg | 95 |
| 42 | Mic | Cell | Cancer Cell Det/Seg | 96 |
| 43 | US | Chest | Cancer Det/Seg | 97 |
| 44 | US | Thyroid | Thyroid Nodule Region Det/Seg | 98 |
| 45 | MRI | Intestine | Multi-labels Det/Seg | 103 |
| 46 | MRI | Liver | Liver Det/Seg | 104, 105 |
| 47 | X-ray | Lung | Lung Det/Seg | 99 |
| 48 | X-ray | Lung | Pneumothorax Det/Seg | 106 |
| 49 | X-ray | Bones | Spinal Anomaly Det | 100 |
| 50 | X-ray | Chest | Multi-labels Det | 101, 102 |
| 51 | FP | Fundus | Vessel Seg | 107 |
| 52 | FP | Fundus | Optic Disc and Cup Seg | 108 |
| 53 | FP | Fundus | Optic Disc Seg | 109 |
| No. | Name | Description | Citation |
| 1 | Intel & MobileODT Cervical Screening | Cervix Type in Screening | BenO et al. (2017) |
| 2 | CT Kindney Dataset | Normal or Cyst or Tumor | Islam et al. (2022a) |
| 3 | SARS-COV-2 Ct-Scan | COVID19, Classification Dataset | Soares et al. (2020) |
| 4 | COVID CT COVID-CT | COVID19, Classification Dataset. | Zhao et al. (2020) |
| 5 | Chest CT-Scan | Cancer Classification | SunneYi (2021) |
| 6 | COVID-19-CT SCAN IMAGES | COVID19, Classification | wjXiaochuangw (2019) |
| 7 | Head CT | Head Hemorrhage | Kitamura (2018) |
| 8 | CT of Brain | Head Cancer | Data (2023) |
| 9 | MED-NODE | Melanoma or Naevus | Giotis et al. (2015) |
| 10 | ISIC 2020 | Melanoma, Benign or Malignant | Rotemberg et al. (2021) |
| 11 | PED-UFES-20 | Skin Multi Classification | Pacheco et al. (2020) |
| 12 | Web-scraped Skin Image | Skin Desease Multi Classification | Islam et al. (2022b) |
| 13 | ISBI 2016 | Skin Lesion Classification | Gutman et al. (2016) |
| 14 | ISIC 2019 | Skin Desease Multi Classification | Combalia et al. (2019) |
| 15 | Skin Cancer ISIC | Skin Cancer Multi Classification | Katanskiy (2019) |
| 16 | Dental Condition Dataset | Teeth condition classification | Sajid (2024) |
| 17 | Oral Cancer Dataset | Oral cancer Classification | RASHID (2024) |
| 18 | The Nerthus Dataset | Cleanliness level | Pogorelov et al. (2017a) |
| 19 | Endoscopic Bladder Tissue | Canser Degree Classification | Lazo et al. (2023) |
| 20 | Kvasir | Multi Disease Classification | Pogorelov et al. (2017b) |
| 21 | ACRIMA | Glaucoma | Ovreiu et al. (2021) |
| 22 | Augemnted ocular diseases AOD | Multi Classification of eye diseases | Бақтыбекұлы (2021) |
| 23 | JSIEC | Multi Classification of eye diseases | Cen et al. (2021) |
| 24 | Multi-Label Retinal Diseases | Multi Classification of eye diseases | Rodríguez et al. (2022) |
| 25 | RFMiD 2.0 | Multi Classification of eye diseases | Panchal et al. (2023) |
| 26 | ToxoFundus(Data Processed Paper) | Ocular toxoplasmosis | Cardozo et al. (2023) |
| 27 | ToxoFundus(Data Raw 6class All) | Ocular toxoplasmosis | Cardozo et al. (2023) |
| 28 | Adam dataset | Age-related Macular Degeneration | Liang (2021) |
| 29 | APTOS 2019 Blindness | Blindness Level Identification 0 4 | Karthik et al. (2019) |
| 30 | DRIMBD | Quality Testing of Retinal Images | Prentasic et al. (2013) |
| 31 | Glaucoma Detection | Glaucoma Classification | Zhang and Das (2022) |
| 32 | AIROGS | Glaucoma Classification | de Vente et al. (2023) |
| 33 | ICPR-HEp-2 | Multi Classification | Qi et al. (2016) |
| 34 | SICAPv2 | Cancer Degree Classification | Silva-Rodríguez et al. (2020) |
| 35 | Blood Cell Images | Blood Cell Classificaion (Multi) | Mooney (2017) |
| 36 | BreakHis | Cell type and beginormag | Bukun (2019) |
| 37 | Chaoyang | Multi Classification of pathologists | Zhu et al. |
| 38 | HuSHeM | Sperm Head Morphology Classificaion | Shaker (2018) |
| 39 | Bone Marrow Cell Classification | Bone Marrow Cell Classification | Matek et al. (2021) |
| 40 | NCT-CRC-HE-100K | Multi Classification | Kather et al. (2018) |
| 41 | Malignant Lymphoma Classification | Multi Classification | Orlov et al. (2010a) |
| 42 | Histopathologic Cancer Detection | Cancer Classification | Cukierski (2018) |
| 43 | LC25000 | Multi Classification of Lung and Colon | Zhu (2022) |
| 44 | Brain Tumor 17 Classes | Multi Classification | Feltrin (2022) |
| 45 | Tumor Classification | Pituitary or Glioma or Meningioma or Notumor | Nickparvar (2021a) |
| 46 | Malignant Lymphoma Classification | Multi Classification of eye diseases | Orlov et al. (2010b) |
| 47 | Retinal OCT-C8 | Multi Classification of eye diseases | Subramanian et al. (2022) |
| 48 | BUSI | Breast Cancer | Al-Dhabyani et al. (2020) |
| 49 | Digital Knee X-Ray Images | Degree Classification of Knee | Gornale and Patravali (2020) |
| 50 | Bone Fracture Multi-Region X-ray Data | Fractured Classification | Nickparvar (2021b) |
| 51 | Fracture detection | Fractured Classification | Batra (2024) |
| 52 | The vertebrae X-ray image | Vertebrae | Fraiwan et al. (2022) |
| 53 | Knee Osteoarthritis Dataset | Knee Osteoarthritis with severity grading | Chen (2018) |
| 54 | Shenzhen Chest X-Ray Set | COVID19, Classification Dataset. | Jaeger et al. (2014) |
| 55 | Chest X-ray PD | COVID and Pneumonia | Asraf and Islam (2021) |
| 56 | COVID-19 CHEST X-RAY DATABASE | COVID and Pneumonia | Chowdhury et al. (2020) |
| 57 | COVIDGR | COVID19, Classification | Tabik et al. (2020) |
| 58 | MIAS | Multi Classification of Breast | Mader (2017) |
| 59 | Tuberculosis Chest X-Ray Database | Tuberculosis | Rahman et al. (2020) |
| 60 | Pediatric Pneumonia Chest X-Ray | Pneumonia Classification | Kermany (2018) |
| No. | Name | Description | Citation |
| 61 | Random Sample of NIH Chest X-Ray Dataset | Multi Classificaiton of Chest | Wang et al. (2017) |
| 62 | CoronaHack-Chest X-Ray | Pnemonia Classifcition with Virus type | Praveen (2019) |
| 63 | Brain Tumor Dataset | Tumor Classification | Viradiya (2020) |
| 64 | Fitzpatrick 17k (Nine Labels) | Multi Classification | Groh et al. (2021) |
| 65 | BioMediTech | Multi Classification | Nanni et al. (2016) |
| 66 | Diabetic retinopathy | Diabetic Retinopathy Level | Benítez et al. (2021) |
| 67 | Leukemia | Cancer Classification | Codella et al. (2019) |
| 68 | ODIR-5K | Multiple Labels Classification | University (2019) |
| 69 | Arthrosis | Bone Age Classification | Zha (2021) |
| 70 | HSA-NRL | Multi Classification of pathologists | Zhu et al. (2021) |
| 71 | ISIC 2018 (Task 3) | Multi Classification | Codella et al. (2019) |
| 72 | ISIC 2017 (Task 3) | Multi Classification | Codella et al. (2018) |
| 73 | ChestX-Det | Multi Classification | Lian et al. (2021) |
| 74 | Monkeypox Skin Lesion Dataset | Only Monkeypox | Ali et al. (2022) |
| 75 | Cataract Dataset | Multi Classification | JR2NGB (2019) |
| 76 | ChestX-rays IndianaUniversity | Multi-label Classification | Raddar (2019) |
| 77 | CheXpert v1.0 small | Multi-label Classification | Arevalo (2020) |
| 78 | CBIS-DDSM | Multi Classification | Lee et al. (2017) |
| 79 | NLM-TB | Tuberculosis | Karaca (2022) |
| 80 | ChestXray-NIHCC | Multi-label Classification | Summers and Ronald (2020) |
| 81 | COVIDx CXR-4 | COVID19, Classification | Wang et al. (2020) |
| 82 | VinDr-Mammo | Multi-label Classification | Nguyen et al. (2023) |
| 83 | PBC dataset normal DIB | Multi Classification | Acevedo et al. (2020) |
| 84 | Human Protein Atlas | Multi-label Classification (Only green) | Le et al. (2022) |
| 85 | RSNA Pneumonia Detection Challenge 2018 | Multi-label Classification | Anouk Stein et al. (2018) |
| 86 | VinDr-SpineXR | Multi Classification of Bones Diseases | Pham et al. (2021) |
| 87 | VinDr-PCXR | Multi-label Classification | Pham et al. (2022) |
| 88 | PH2 | Melanoma Segmentation | Mendonca et al. (2015) |
| 89 | ISBI 2016 (Task3B) | Melanoma Segmentation | Gutman et al. (2016) |
| 90 | ISIC 2016 (Task 1) | Melanoma Segmentation | Gutman et al. (2016) |
| 91 | ISIC 2017 | Melanoma Segmentation | Codella et al. (2018) |
| 92 | CVC-ClinicDB | Polyp Segmentation | Bernal et al. (2015) |
| 93 | Kvasir-SEG | Polyp segmentation | Jha et al. (2020) |
| 94 | m2caiseg | Surgical Instrument Segmentation | Maqbool et al. (2020) |
| 95 | EDD 2020 | Multiple Diseases Segmentation in Intestine | Ali et al. (2020) |
| 96 | SICAPv2 | Cancer Cells Segmentation | Silva-Rodríguez et al. (2020) |
| 97 | BUSI | Cancer Segmentation | Hesaraki (2022) |
| 98 | TN3K | Thyroid Nodule Segmentation | Gong et al. (2022) |
| 99 | NLM-TB | Lung Segmentation (With left or right) | Gong et al. (2021) |
| 100 | VinDr-SpineXR | Spinal X-ray Anaomaly Detection | Pham et al. (2021) |
| 101 | VinDr-PCXR | Multiple Diseases Segmentation in Chest | Pham et al. (2022) |
| 102 | ChestX-Det | Multiple Diseases Segmentation in Chest | Lian et al. (2021) |
| 103 | UW-Madison Gl Tract Image Segmentation | Surgical Instrument Segmentation | Lee et al. (2024) |
| 104 | Duke Liver Dataset MRI v1 | Liver Segmentation | Macdonald et al. (2020) |
| 105 | Duke Liver Dataset MRI v2 | Liver Segmentation | Macdonald et al. (2020) |
| 106 | SIIM-ACR Pneumothorax Segmentation | Pneumothorax Segmentation | Zawacki et al. (2019) |
| 107 | FIVES | Fundus Vascular Segmentation | Jin et al. (2022) |
| 108 | RIM-ONE DL | Optic Disc and Cup Segmentation | Batista et al. (2020) |
| 109 | PALM19 | Optic Disc Segmentation | Fu et al. (2019) |

### v2

@@
-On the Compositional Generalization of Multimodal LLMs for Medical Imaging
+Exploring Compositional Generalization of Multimodal LLMs for Medical Imaging
 Abstract
-Multimodal large language models (MLLMs) hold significant potential in the medical field, but their capabilities are often limited by insufficient data in certain medical domains, highlighting the need for understanding what kinds of images can be used by MLLMs for generalization.
-Current research suggests that multi-task training outperforms single-task as different tasks can benefit each other, but they often overlook the internal relationships within these tasks, providing limited guidance on selecting datasets to enhance specific tasks.
-To analyze this phenomenon, we attempted to employ compositional generalization (CG)—the ability of models to understand novel combinations by recombining learned elements—as a guiding framework.
-Since medical images can be precisely defined by Modality, Anatomical area, and Task, naturally providing an environment for exploring CG.
-Therefore, we assembled 106 medical datasets to create Med-MAT for comprehensive experiments.
+Medical imaging provides essential visual insights for diagnosis, and multimodal large language models (MLLMs) are increasingly utilized for its analysis due to their strong generalization capabilities; however, the underlying factors driving this generalization remain unclear.
+Current research suggests that multi-task training outperforms single-task as different tasks can benefit each other, but they often overlook the internal relationships within these tasks.
+To analyze this phenomenon, we attempted to employ compositional generalization (CG), which refers to the models’ ability to understand novel combinations by recombining learned elements, as a guiding framework.
+Since medical images can be precisely defined by Modality, Anatomical area, and Task, naturally providing an environment for exploring CG, we assembled 106 medical datasets to create Med-MAT for comprehensive experiments.
 The experiments confirmed that MLLMs can use CG to understand unseen medical images and identified CG as one of the main drivers of the generalization observed in multi-task training.
-Additionally, further studies demonstrated that CG effectively supports datasets with limited data and delivers consistent performance across different backbones, highlighting its versatility and broad applicability.
-Med-MAT is publicly available at github.com/FreedomIntelligence/Med-MAT.
+Additionally, further studies demonstrated that CG effectively supports datasets with limited data and confirmed that MLLMs can achieve CG across classification and detection tasks, underscoring its broader generalization potential.
+Med-MAT is available at https://
+github.com/FreedomIntelligence/Med-MAT.
 UTF8gkai
-On the Compositional Generalization of Multimodal LLMs
+Exploring Compositional Generalization of Multimodal LLMs
 for Medical Imaging
@@
 1 Introduction
-Multimodal large language models (MLLMs) Liu et al.
+Medical imaging provides essential visual insights into the structures of the human body, making it a critical tool for medical diagnosis.
+Recently, multimodal large language models (MLLMs) Liu et al.
 (2023); Li et al.
 (2024); Chen et al.
-(2024b) are showing great promise for the medical community, facilitating efficient consultations for doctors and providing patients with anytime access to their medical conditions.
-However, limited data on rare or privacy-restricted diseases often restricts MLLMs’ performance, making it crucial to explore what kinds of images can be used by MLLMs to generalize these diseases or learn them efficiently.
-Current research on medical image generalization Mo and Liang (2024); Ren et al.
+(2024b) have been employed to analyze these images due to their strong interpretability and generalization capabilities.
+In this paper, we focus on the latter: generalization of MLLMs in medical imaging.
+Current research Mo and Liang (2024); Ren et al.
 (2024) has demonstrated that models trained on multiple tasks outperform those trained on a single task as they can leverage potential knowledge from other tasks.
-However, these studies have not conducted a detailed analysis of which data within multitask datasets can mutually complement each other.
-Consequently, we aim to investigate the phenomenon of mutual improvement in MLLMs’ understanding of medical images from the perspective of composition generalization.
-Compositional generalization (CG) Xu et al.
-(2022); Li et al.
-(2019) refers to a model’s ability to learn fundamental elements and recombine them in novel ways to understand unseen combinations (Figure 1).
-In medical imaging, each image can be categorized by Modality, Anatomical area, and medical Task, presenting numerous natural opportunities for CG.
+Yet, the underlying factors that contribute to this generalization remain insufficiently explored.
+To this end, we take the perspective of composition generalization (CG) Li et al.
+(2019); Xu et al.
+(2022); Tang et al.
+(2024) to investigate the generalization phenomenon of mutual improvement in MLLMs’ understanding of medical images.
+Specifically, CG is the model’s ability to learn fundamental elements and recombine them in novel ways to understand unseen combinations (e.g., learning Cat from White Cat and Black from Black Dog, then generalizing to Black Cat, as shown in Figure 1).
+In this paper, we categorize each image to three elements: Modality
+, Anatomical area
+, and medical Task
+, presenting numerous natural opportunities for CG.
 We defined these three elements as the MAT-Triplet and collected 106 medical datasets, subsequently merging those that share the same MAT-Triplet to create the Med-MAT dataset.
 Ultimately, Med-MAT comprises 53 subsets, encompassing 11 modalities, 14 anatomical regions, and 13 medical tasks, providing a foundation for investigating CG and other generalization methods.
-To verify the existence of CG, we designated certain datasets as Target data and selected all the Related data from Med-MAT that shared the same MAT-Triplet with the Target data for generalization.
-In addition, we conducted larger-scale experiments and observed the changes in model generalization performance after deliberately disrupting CG.
-Ultimately, the experiments demonstrated that MLLMs can leverage CG to understand unseen medical images and CG is one of the key factors driving the generalization observed in multi-task training.
-Subsequently, the practical applications of CG and its effects on different backbones were further explored.
-We found that increasing the volume of CG combinations continuously enhanced the model’s understanding of the Target data.
-While CG’s direct impact on unseen data was not always apparent, CG could assist the model in efficient fitting with limited data by being incorporated into the training data of the Target.
-Furthermore, we discovered that CG exists in different MLLM backbones, underscoring CG’s broad applicability.
+To verify the existence of CG, we designated specific datasets as Target data and selected all Related data from Med-MAT that shared the same MAT-Triplet with the Target data.
+Using these data combinations, we accessed the generalization performance of MLLMs and observed that they could leverage CG to understand unseen medical images.
+To further validate this finding, we repeated the experiments on different MLLMs and obtained consistent results, confirming the universality of CG.
+Building on these insights, we expanded the number of combinations and observed the changes in model generalization performance after deliberately disrupting CG, ultimately revealing that CG is a key factor driving the generalization of MLLMs.
+Furthermore, we explored the potential applications of CG and its performance across classification and detection tasks, finding that CG enhances MLLMs’ ability to handle medical scenarios with limited training data and improves their capacity for spatial awareness.
 Here are the key contributions of our work: 1) A VQA dataset, Med-MAT, has been constructed, providing a platform to explore the generalization of MLLMs on medical images.
-2) Through this dataset, we observed that MLLMs can utilize compositional generalization to understand unseen images, demonstrating that this is one of the main forms of generalization for medical MLLMs.
-3) Finally, the contribution of CG to understanding medical domains with limited data has been further explored, with its presence across different backbones demonstrating its broad applicability.
-| Model | 02 | 03 | 07 | 08 | 09 | 11 | 13 | 14 | 15 | 16 | 18 | 19 | 21 | 22 | 23 | 25 | 26 | 28 | 30 | 31 | 32 | 33 | 35 | 36 | 37 |
-| Baseline | 22 | 47 | 40 | 25 | 26 | 27 | 28 | 24 | 22 | 24 | 25 | 23 | 49 | 26 | 25 | 24 | 49 | 30 | 49 | 21 | 49 | 20 | 25 | 23 | 19 |
+2) Through this dataset, we observed that MLLMs in different architectures can utilize compositional generalization to understand unseen images and demonstrated that this is one of the main forms of generalization for medical MLLMs.
+3) Finally, the real-world applicability of CG, along with its presence across detection and classification tasks, has been further explored, highlighting its potential to enhance data-efficient training and its broad applicability.
+| Subset No. | 02 | 03 | 07 | 08 | 09 | 11 | 13 | 14 | 15 | 16 | 18 | 19 | 21 | 22 | 23 | 25 | 26 | 28 | 30 | 31 | 32 | 33 | 35 | 36 | 37 |
+| Baseline | 21 | 47 | 40 | 25 | 26 | 27 | 28 | 24 | 22 | 24 | 25 | 23 | 49 | 26 | 25 | 24 | 49 | 30 | 49 | 21 | 49 | 20 | 25 | 23 | 19 |
 | Single-task Training | 24 | 49 | 50 | 68 | 65 | 76 | 83 | 53 | 61 | 32 | 29 | 26 | 57 | 53 | 28 | 24 | 57 | 64 | 89 | 60 | 97 | 54 | 29 | 51 | 49 |
 | Multi-task Training | 96 | 89 | 80 | 80 | 79 | 97 | 92 | 88 | 76 | 57 | 88 | 74 | 87 | 86 | 93 | 52 | 98 | 72 | 94 | 61 | 100 | 72 | 75 | 60 | 50 |
-| Model | 01 | 04 | 05 | 06 | 10 | 12 | 17 | 20 | 24 | 27 | 29 | 34 |
+| Subset No. | 01 | 04 | 05 | 06 | 10 | 12 | 17 | 20 | 24 | 27 | 29 | 34 |
 | Baseline | 32 | 25 | 33 | 33 | 48 | 27 | 33 | 13 | 34 | 37 | 31 | 20 |
 | Multi-task Training | 39 | 26 | 70 | 31 | 58 | 38 | 61 | 40 | 35 | 41 | 55 | 50 |
-2 Med-MAT
+2 A Pilot Study on Generalization
+2.1 Data Collection (Med-MAT)
 Most existing datasets for MLLMs Zhang et al.
@@
 To address this gap, we curated a large collection of image-text pairs to develop Med-MAT, ensuring that each sample is explicitly defined by MAT-Triplet.
-Section 2.1 provides data processing methods; and Section 2.2 presents a pilot experiment.
-2.1 Data Processing
-Data Construction: Med-MAT contains a total of 106 image-label pair medical datasets, sourced from various medical public challenges or high-quality annotated datasets.
+Data Construction
+Med-MAT contains a total of 106 image-label pair medical datasets, sourced from various medical public challenges or high-quality annotated datasets.
 All datasets are categorized according to their MAT-Triplet, with data having identical elements grouped into a single subset (Figure 2).
 Labels are manually clustered to ensure that annotations with the same meaning are not repeatedly used.
-In total, Med-MAT covers 11 medical modalities, 14 anatomical areas, and 13 medical tasks, hoping that it can spread across various medical tasks like a mat.
-(Data lists are shown in Appendix B)
-Data Distribution: All subsets are divided into training and test sets following their original distributions or using a 9:1 ratio.
-To ensure a fair comparison, each training set is limited to 3,000 samples, with label balance maintained as much as possible.
+In total, Med-MAT covers 11 medical modalities
+, 14 anatomical areas
+, and 13 medical tasks
+, hoping that it can spread across various medical tasks like a mat. (Data lists are shown in Appendix B)
+Data Distribution
+All subsets are divided into training and test sets following their original distributions or using a 9:1 ratio.
+To ensure a fair comparison, each training set is limited to 3,000 samples 111Most datasets contain around 3,000 samples., with label balance maintained as much as possible.
 Any subset that cannot meet this requirement is treated as an OOD (out-of-distribution) dataset.
 For the test sets, we strictly balance the number of samples per label to ensure that the accuracy metric reliably reflects model performance.
-QA Pairs Construction: To enable MLLMs to directly train and test on Med-MAT, all image-label paired data were converted into a visual question-answering (VQA) format (Figure 3).
+QA Pairs Construction
+To enable MLLMs to directly train and test on Med-MAT, all image-label paired data were converted into a visual question-answering (VQA) format (Figure 3).
 Specifically, each subset was manually assigned 6 instructions to guide the MLLM in answering the subset task.
@@
 (2024b), a non-medical dataset consisting of single-answer, four-option questions, was incorporated during the training.
-2.2 A Pilot Study of Data Composition
-Experiments Setup: Experiments in this paper mainly focused on classification datasets to explore the model’s combinatorial generalization capabilities across various image types.
-The base model LLaVA-v1.5-7B-Vicuna Liu et al.
-(2023) features a transparent pretraining process and uses very little medical data during training, thereby minimizing the risk of knowledge leakage.
-By leveraging the flexibility of MLLM, we achieved task switching and generalization simply by modifying the corresponding prompts, significantly simplifying generalization studies.
-Each experiment was conducted over 5 epochs using 8 A800 (80GB) GPUs, with a batch size of 32 and a learning rate of 5e-6.
-Generalization performance was assessed by comparing the model’s accuracy on the Target data.
-Pilot Study: To explore the benefits of data composition for downstream tasks, all in-distribution (ID) datasets were combined for multi-task training, with single-task training on individual ID datasets serving as the control.
-The results (Table 1 and 2) showed that multi-task training outperformed single-task training on specific tasks, and also surpassed the baseline in predicting unseen OOD datasets.
-This suggests that certain data combinations can enhance the model’s classification performance, highlighting the potential for CG.
-Exploring which combinations provide valuable insights for downstream medical tasks in a promising avenue for further research.
-Take-away 1: Certain data combinations can assist MLLMs in medical classification tasks.
+2.2 Observation
+Experiment Setup We chose LLaVA-v1.5-7B-Vicuna Liu et al.
+(2023) as the base model due to its transparent pretraining process and minimal use of medical data, reducing the risk of knowledge leakage.
+Leveraging MLLM’s flexibility, we enabled task switching and generalization by adjusting prompts, streamlining generalization studies.
+Each experiment ran for 5 epochs on 8 A800 GPUs with a batch size of 32 and a learning rate of 5e-6.
+Analysis
+To access the generalization of MLLMs, we trained the baseline on all ID datasets to simulate Multi-task Training and separately trained on individual ID datasets to establish the Single-task Training as the control group.
+We then evaluated the models on all datasets.
+The results in Table 1 and 2 confirm that Multi-task Training outperformed Single-task Training on specific tasks and improved OOD prediction, suggesting certain data combinations enhance classification and identifying valuable combinations for medical tasks warrants further research.
+This observation leads to a research question (RQ):
+What drives the generalization observed in MLLMs during Multi-task Training?
+To address it, we aim to explore the generalization mechanism of MLLMs from the perspective of compositional generalization (CG).
 3 Proof of Concept on CG
-To explore whether MLLMs can leverage CG to understand unseen images, controlled variable studies were first conducted using all possible CG combinations to determine if CG exists among the MAT-Triplet (Section 3.1).
-Subsequently, the number of combinations was increased to explore more generalizable insights, aiming to assess the performance gains of multi-task training (Section 3.2).
-| Related Combination | Target Subset | Baseline | Baseline+ | Trained | |
-| Lung, COVID | Brain, Cancer | Lung, Cancer | 25 | 25 | 27 |
-| Lung, Cancer | Brain, State | Lung, State | 47 | 46 | 50 |
-| Brain, Cancer | Lung, State | Brain, State | 33 | 50 | 57 |
-| Bones, Level | Lung, State | Bones, State | 49 | 53 | 51 |
-| Bones, Level | Brain, State | Bones, State | 49 | 53 | 72 |
-| Bones, Level | Breast, Diseases | Bones, Diseases | 37 | 33 | 39 |
-| Bones, Level | Lung, Diseases | Bones, Diseases | 37 | 33 | 43 |
-| Bones, Level | Chest, Diseases | Bones, Diseases | 37 | 31 | 43 |
-| Bones, State | Breast, Diseases | Bones, Diseases | 37 | 37 | 43 |
-| Bones, State | Lung, Diseases | Bones, Diseases | 37 | 37 | 43 |
-| Bones, State | Chest, Diseases | Bones, Diseases | 37 | 37 | 41 |
-| Lung, COVID | Breast, Diseases | Lung, Diseases | 49 | 48 | 51 |
-| Lung, COVID | Bones, Diseases | Lung, Diseases | 49 | 48 | 52 |
-| Lung, COVID | Chest, Diseases | Lung, Diseases | 49 | 48 | 51 |
-| CT, Cancer | X-ray, COVID | CT, COVID | 47 | 46 | 72 |
-| CT, COVID | X-ray, Diseases | X-ray, COVID | 30 | 21 | 49 |
-| CT, State | X-ray, Diseases | X-ray, State | 30 | 21 | 46 |
-| CT, State | X-ray, Cancer | CT, Cancer | 33 | 28 | 28 |
-| CT, Brain(State) | X-ray, Bones | X-ray, Brain | 49 | 49 | 91 |
-| CT, Brain | X-ray, Lung | X-ray, Brain | 49 | 50 | 81 |
-| CT, Brain(Cancer) | X-ray, Bones | X-ray, Brain | 25 | 51 | 74 |
-| CT, Brain | X-ray, Lung | X-ray, Brain | 49 | 52 | 52 |
-| X-ray, Brain | CT, Lung(State) | CT, Brain(State) | 33 | 50 | 60 |
-| X-ray, Lung | CT, Brain | CT, Lung(Cancer) | 25 | 25 | 36 |
-| X-ray, Lung | CT, Brain(State) | CT, Lung | 47 | 50 | 81 |
-| X-ray, Lung | CT, Brain(Cancer) | CT, Lung | 47 | 50 | 71 |
-| CT, Lung (State) | X-ray, Bones | X-ray, Lung | 30 | 32 | 28 |
-| CT, Lung (State) | X-ray, Brain | X-ray, Lung | 30 | 32 | 35 |
-| CT, Lung (Cancer) | X-ray, Bones | X-ray, Lung | 30 | 32 | 41 |
-| CT, Lung (Cancer) | X-ray, Brain | X-ray, Lung | 30 | 32 | 42 |
-| Der, Skin, Cancer | FP, Fundus, Diseases | Der, Skin, Diseases | 25 | 29 | 33 |
-| Der, Skin, Cancer | OCT, Retine, Diseases | Der, Skin, Diseases | 25 | 29 | 33 |
-| Der, Skin, Diseases | DP, Mouth, Cancer | Der, Skin, Cancer | 40 | 33 | 63 |
-| Der, Skin, Diseases | Mic, Cell, Cancer | Der, Skin, Cancer | 40 | 33 | 63 |
-| DP, Mouth, State | Der, Skin, Cancer | DP, Mouth, Cancer | 48 | 50 | 52 |
-| DP, Mouth, State | Mic, Cell, Cancer | DP, Mouth, Cancer | 48 | 50 | 55 |
-| FP, Fundus, Diseases | Mic, Cell, Level | FP, Fundus, Level | 33 | 36 | 42 |
-| Mic, Cell, Cell Identification | FP, Fundus, Level | Mic, Cell, Level | 23 | 33 | 32 |
-| Mic, Cell, Cell identification | Der, Skin, Cancer | Mic, Cell, Cancer | 49 | 50 | 50 |
-| Mic, Cell, Cell identification | DP, Mouth, Cancer | Mic, Cell, Cancer | 49 | 51 | 62 |
-| Mic, Cell, Level | Der, Skin, Cancer | Mic, Cell, Cancer | 49 | 51 | 52 |
-| Mic, Cell, Level | DP, Mouth, Cancer | Mic, Cell, Cancer | 49 | 51 | 58 |
-| Mic, Cell, Cancer | FP, Fundus, Level | Mic, Cell, Level | 23 | 24 | 27 |
-3.1 Controlled Variable Study on CG
-3.1.1 Experiment Setup
-In this section, the existence of CG was explored from a finer perspective, focusing on CG with only two elements varying while the third remained constant.
-For example, we fixed the Modality and examined the CG between Areas and Tasks.
-Additionally, we identified specific Modality-Area pairs, such as dermoscopy paired consistently with skin, which were treated as a special category.
-These 4 different fixed formats were classified into distinct Direction Types.
-Following the setup in Section 2.2, we evaluated the model’s performance on Target data.
+This section will prove the existence of CG in MLLMs, offering preliminary insights to address the RQ and providing support for further analysis.
+| Related Combination | Target Subset | Baseline | Baseline+ | Trained | CG Helps | ||||
+25 | 25 | 27 | ✓ |
+47 | 46 | 50 | ✓ |
+33 | 50 | 57 | ✓ |
+49 | 53 | 51 | ✗ |
+49 | 53 | 72 | ✓ |
+37 | 33 | 39 | ✓ |
+37 | 33 | 43 | ✓ |
+37 | 31 | 43 | ✓ |
+37 | 37 | 43 | ✓ |
+37 | 37 | 43 | ✓ |
+37 | 37 | 41 | ✓ |
+49 | 48 | 51 | ✓ |
+49 | 48 | 52 | ✓ |
+49 | 48 | 51 | ✓ |
+47 | 46 | 72 | ✓ |
+30 | 21 | 49 | ✓ |
+30 | 21 | 46 | ✓ |
+33 | 28 | 28 | ✗ |
+49 | 49 | 91 | ✓ |
+49 | 50 | 81 | ✓ |
+25 | 51 | 74 | ✓ |
+49 | 52 | 52 | ✗ |
+33 | 50 | 60 | ✓ |
+25 | 25 | 36 | ✓ |
+47 | 50 | 81 | ✓ |
+47 | 50 | 71 | ✓ |
+30 | 32 | 28 | ✗ |
+30 | 32 | 35 | ✓ |
+30 | 32 | 41 | ✓ |
+30 | 32 | 42 | ✓ |
+25 | 29 | 33 | ✓ |
+25 | 29 | 33 | ✓ |
+40 | 33 | 63 | ✓ |
+40 | 33 | 63 | ✓ |
+48 | 50 | 52 | ✓ |
+48 | 50 | 55 | ✓ |
+33 | 36 | 42 | ✓ |
+23 | 33 | 32 | ✗ |
+49 | 50 | 50 | ✗ |
+49 | 51 | 62 | ✓ |
+49 | 51 | 52 | ✓ |
+49 | 51 | 58 | ✓ |
+23 | 24 | 27 | ✓ |
+3.1 Experiment Setup
+To explore the existence of CG from a finer perspective, this section focuses on CG with only two MAT-Triplet elements varying while the third remains constant.
+Additionally, we identified specific Modality-Area pairs
+, such as dermoscopy paired consistently with skin, which were treated as a special category. These 4 different fixed formats were classified into distinct Direction Types.
+We adhered to the training setup described in Section 2.2 and evaluated the model’s performance on the Target data.
 Baseline refers to the model without any training, while Trained refers to the model trained solely on Related data.
 To ensure that our conclusions are not influenced by the amount of training data, we randomly sampled an equal number of data from the Unrelated subsets, and this configuration is referred to as Baseline+.
-3.1.2 Analysis of CG in MLLMs
+3.2 Results
 Results are shown in Table 3 and it can be observed that almost all CG combinations are able to generalize to downstream tasks, highlighting that MLLMs can leverage CG to generalize Target data across all Direction Types.
-Besides that, since this experiment focused solely on two-element tuples, we further investigated three-element tuples in Section 4.3, where we also observed similarly strong generalizations when obtaining MAT-Triplet elements from three different datasets.
-Take-away 2: MLLMs can leverage CG to understand unseen medical images.
+Besides that, since this experiment focused solely on two-element tuples, we further investigated three-element tuples in Appendix A.4, where we also observed similarly strong generalizations when obtaining MAT-Triplet elements from three different datasets.
+Take-away 1: MLLMs can leverage CG to understand unseen medical images.
+| Related Combination | Target Subset | Qwen | Llama | ||||
 In the Baseline+ setting, we removed all datasets sharing any MAT-Triplet element with the Target data.
@@
 This suggests that only datasets related through the MAT-Triplet can help the model learn and generalize to new target tasks.
-Take-away 3: Generalization only arises from MAT-Triplet related medical datasets.
-3.1.3 Analysis of Non-obvious Cases
-Some Trained models show minimal gains or performance declines, especially in Level or Diseases Identification tasks.
-The images in the Level Identification are very similar, while the Diseases Identification involves a large number of target labels, both of which require large datasets for training, which may explain the poor performance.
-To further analyze them, we modified the evaluation scenario to use CG combinations for assisting in Target data training with limited samples (Section 4.2).
-3.2 Scaling the Combination Number of CG
-To ensure the universality of CG, we expanded the number of combination datasets to evaluate its effects on a larger scale.
-This section explores the following questions: (Q1) While Meta CG experiments indicate that Unrelated combinations provide no benefit to Target data, can generalization arise when training incorporates more Unrelated combinations, simulating a multi-task scenario?
+Take-away 2: Generalization arises in medical datasets in which at least partial MAT elements pre-exist during training.
+3.3 Extending CG to other Backbones
+LLaVA was selected as the baseline because its training data and processes are publicly available, ensuring minimal exposure to medical images and preventing bias in the integration of medical image knowledge into the MLLM.
+To ensure that the results are not affected by the training data or the visual encoder of LLaVA, we randomly sampled two combinations from each Direction Type to investigate CG on Qwen2-VL-7B Wang et al.
+(2024a) and Llama3.2-11B-Vision Meta AI (2024).
+Qwen2-VL undergoes additional training on proprietary data based on ViT and incorporates a strategy to adjust the number of vision tokens according to resolution.
+Llama3.2-Vision, on the other hand, pretrains its own vision encoder from scratch using proprietary data.
+Thus, both models serve as a means to assess whether MLLMs with different training data and vision encoders can still leverage CG to understand unseen images, ensuring that CG is not merely an artifact of LLaVA’s data fitting or specific to its vision encoder.
+Table 4 presents the experimental results, showing that both selected backbones exhibit a certain degree of generalization across most tasks.
+This suggests that despite differences in pre-train data and vision encoders, different MLLMs can still leverage CG to understand unseen images.
+Take-away 3: CG persists across different MLLM backbones.
+4 Scaling Combination in CG
+After confirming that CG is indeed a form of generalization in MLLMs, we expanded the number of participating combinations to explore the generalizability of CG and examine its relationship with the generalization exhibited by Multi-task Training to address the RQ.
+4.1 Experiment Setup
+Two sub-questions have been defined to verify the applicability of CG in multiple data combinations and examine its role in Multi-task Training.
+(Q1) While previous experiment on CG indicated that Unrelated combinations provide no benefit to Target data, can generalization arise when training incorporates more Unrelated combinations, simulating a multi-task scenario?
 (Q2) Previous studies suggest that Multi-task Training generally promotes better generalization than single-task training.
-To what extent does this generalization originate from CG?
-3.2.1 Experiment Setup
-Selection Strategy: To ensure a balanced evaluation of Related and Unrelated combinations, Subset 03 and Subset 28 were chosen as Target datasets because they exhibit the most balanced ratios of Related to Unrelated subsets (13:11 for Subset 03 and 11:13 for Subset 28), making them ideal for providing a diverse range of compositions in the scale-up experiments.
+If the CG conditions in Multi-task Training are deliberately disrupted, will the resulting generalization effect be affected?
+Selection Strategy
+To ensure a balanced evaluation of Related and Unrelated combinations, Subset 03 and Subset 28 were chosen as Target datasets because they exhibit the most balanced ratios of Related to Unrelated subsets (13:11 for Subset 03 and 11:13 for Subset 28), making them ideal for providing a diverse range of compositions in the scale-up experiments.
 The baseline was trained on all subsets excluding the Target data to evaluate the claim that mixing multi-task data enhances generalization (All Data).
@@
 To ensure consistency, the total data volume in all experiments was limited to 15,000 samples, aligning with the number of ID subsets available after excluding related tasks from Subset 03.
-3.2.2 Analysis of Scaling Experiment
+4.2 Analysis of Scaling Experiment
 Figure 4 illustrates the results.
@@
 Besides, w/o Modality / Area / Task showed significant accuracy drops compared to All Related, despite holding the training data volume constant.
-Notably, All Related performs at a level comparable to All Data, where all datasets are included in training.
-This suggests that CG provides significant help for the generalization brought by multi-task training.
-Take-away 5: CG is one of the primary forms of generalization for MLLMs trained on multi-task dataset in medical imaging (Q2).
-| Related Combination | Target Subset | Baseline | Trained | ||
-| CT - Subset02 | Brain - Subset22 | Cancer - Subset07 | CT, Brain, Cancer | 28 | 26 |
-| CT - Subset03 | Brain - Subset22 | Cancer - Subset21 | CT, Brain, Cancer | 28 | 25 |
-| CT - Subset02 | Brain - Subset22 | State - Subset09 | CT, Brain, State | 33 | 64 |
-| CT - Subset03 | Brain - Subset22 | State - Subset26 | CT, Brain, State | 33 | 70 |
-| X-ray - Subset25 | Lung - Subset03 | Diseases - Subset02 | X-ray, Lung, Diseases | 30 | 45 |
-| X-ray - Subset26 | Lung - Subset03 | Diseases - Subset02 | X-ray, Lung, Diseases | 30 | 38 |
-| X-ray - Subset26 | Lung - Subset03 | Diseases - Subset08 | X-ray, Lung, Diseases | 30 | 44 |
-| X-ray - Subset26 | Breast - Subset24 | Diseases - Subset02 | X-ray, Breast, Diseases | 31 | 32 |
-| X-ray - Subset28 | Breast - Subset24 | Diseases - Subset08 | X-ray, Breast, Diseases | 31 | 52 |
-4 In-depth Analysis of CG
-To further explore the application and applicability of CG, several research questions were proposed in this section.
-In order to present rich results with minimal consumption of computational resources, specific data Selection Strategies were applied for each question.
-Here are the research questions:
-(RQ 1): How does the quantity of Related data affect generalization, and should it be maximized to ensure CG quality?
-(RQ 2): Can CG combinations help the model efficiently fit Target data when only a small amount of Target data is available?
-(RQ 3): Previous studies demonstrated that CG exists between two MAT-Triplet elements, does CG still exist if the three MAT-Triplet elements come from three different datasets?
-(RQ 4): Does CG exist between images from Detection tasks and Classification tasks?
-(RQ 5): Can different MLLM backbones leverage CG to interpret unseen medical images?
-4.1 The Relationship Between CG Performance and Data Volume
-Selection Strategy: This section focuses on examining how the data volume of Related combinations influences generalization.
+This indicates that if the CG combinations are forcibly disrupted, MLLMs will lose a significant amount of generalization capability for the target data.
+Take-away 5: Disrupting CG leads to a significant decline in generalization ability. (Q2).
+Notably, All Related achieves a performance level comparable to All Data, where all datasets are included in training.
+This suggests that CG plays a crucial role in enhancing the generalization effect of Multi-task Training.
+Therefore, in conclusion:
+Take-away 6: CG plays an important role in generalization for MLLMs in medical imaging.
+5 Potential Applications of CG
+As MLLMs can use CG to generalize unseen medical images, this section attempts to explore its potential applications in training medical MLLMs.
+5.1 Generalization without Target Data
+In medical tasks, new and unpredictable conditions, like COVID-19, can emerge at any time.
+Exploring how to use CG to help MLLMs enhance their ability to identify unknown diseases in the absence of specific datasets is both important and meaningful.
+We selected some Target datasets and trained the MLLMs using Related and Unrelated data to observe their generalization to the Target data.
+The generalization trend was assessed by progressively increasing the size of the combination datasets.
+Selection Strategy
 To highlight the generalization trends, the combinations with strong generalization results were selected from the main experiments.
@@
 If multiple combinations meet the criteria, a random seed of 42 was used to determine the selection.
-In this experiment, the amount of combination data was gradually increased (from 0 to 750, 1500, 2250, and 3000), and the trend of the model’s accuracy on Target data after training was observed.
-The experimental results are shown in Figure 5, where the red line represents the accuracy curve gained from Related combinations, and the purple line represents the gain from Unrelated combinations.
-Models trained on Related combinations demonstrate strong performance across all four Target data sets.
-As the data volume of Related combinations increases, the model’s understanding of Target data improves consistently.
-Take-away 6: Generalization effect improves with increasing CG combination’s volume. (RQ 1)
-4.2 Achieving Data-Efficient Training through CG
-Selection Strategy: This experiment evaluates how CG aids training and dataset fitting with limited Target data.
-To highlight performance trends, four combinations with poor generalization (marked in red in Table 3) were selected across the four Direction Types, ensuring both models in the control group started with similar baseline accuracies.
-When multiple poor-performing combinations existed within a Direction Type, the more challenging datasets were chosen based on Table 1 to better observe potential improvements.
-In this experiment, the amount of Target data was progressively increased (from 0 to 500, 1000, 1500, and 2000), and a fixed number of combination data were incorporated during training.
-Figure 6 displays the results, with red lines showing accuracy gains from Related combination and purple lines from Unrelated ones.
-It can be observed that training with Related combination data reaches peak performance more quickly, demonstrating another form of CG.
-Therefore, although CG performs poorly on these data due to task characteristics, it still helps the model quickly adapt with only a small amount of Target data.
-Take-away 7: Although CG does not always provide direct generalization gains, it can help Target data achieve data-efficient training. (RQ 2)
-4.3 CG with All MAT-Triplet Elements from Different Sources
-In previous controlled experiments (Section 3.1), one element of the MAT-Triplet was kept constant while CG was explored in the remaining two elements.
-To ensure that all the 3 MAT-Triplet elements of the target data originated from three distinct datasets, additional experiments were conducted to further validate the effectiveness of CG.
-For these experiments, all possible combinations meeting the criteria in Med-MAT were selected (Selection Strategy).
-The results presented in Table 4 demonstrate that most combinations can effectively generalize to the Target data.
-Take-away 8: CG remains effective when all elements come from different datasets (RQ 3).
-4.4 Detection-Capable MLLMs can Leverage Detection Data for CG
+Analysis
+The experimental results are shown in Figure 5, where the red line represents the accuracy curve for Related combinations, and the purple line shows the gain from Unrelated combinations.
+The Related combinations group significantly outperformed the Unrelated combinations in terms of generalization across all tasks, with this ability continuing to improve as the data size increased.
+This suggests that Related combinations, leveraging CG, enhance the model’s ability to understand unknown medical tasks.
+Take-away 7: CG might enable MLLMs to handle tasks without dedicated training data.
+5.2 Generalization with Limited Target Data
+This section investigates the benefit of CG for tasks with limited data, e.g. processing medical images in rare conditions.
+Selection Strategy
+To assess generalization in limited data scenarios, we select combinations with poor generalization from Table 3.
+Specifically, for each Direction Type, we randomly choose a CG combination with weak generalization (i.e., rows marked with ✗ in the last column of Table 3).
+For these combinations, we introduce an additional 2,000 examples from the Target data.
+Analysis
+Figure 6 shows the results.
+It can be seen that as we gradually expand the training volume of Target data, adding the Related combination for training enabled the model to reach the peak performance more quickly.
+This suggests that leveraging CG to assist low-data medical scenarios can lead to more data-efficient training, even when CG does not directly result in significant generalization gains in these scenarios.
+Take-away 8: Although CG might not provide direct generalization gains, it helps data efficiency for MLLM training.
+6 CG across Detection and Classification
 Previous studies Ren et al.
 (2024); Wang et al.
-(2025) have demonstrated that jointly training classification and detection tasks can mutually enhance their performance.
-Building on this, we explored CG within this context.
-To ensure comprehensive conclusions, we conducted experiments using Next-Chat and MiniGPT-v2, which represent the two mainstream approaches for MLLMs to perform segmentation tasks.
-The former treats bounding boxes as embeddings and decodes them into coordinates using an additional visual decoder, while the latter processes coordinate points as special text tokens and generates bounding box coordinates directly as output text.
-The final results (Figure 7) show that all CG combinations demonstrated the model’s successful utilization of detection data for CG to the Target data.
-Take-away 9: MLLMs can combine knowledge from detection tasks with classification tasks to improve classification accuracy (RQ 4).
-4.5 CG Exists Across Different MLLM Backbones
-LLaVA was selected as the baseline because its training data and processes are publicly available, ensuring minimal exposure to medical images and preventing bias in the integration of medical image knowledge into the MLLM.
-To verify that CG is not limited to a single framework, we also selected subsets for experiments on Qwen2-VL-7B Wang et al.
-(2024a) and Llama3.2-11B-Vision Meta AI (2024).
-The results in Figure 7 demonstrate CG effects similar to those observed in the LLaVA.
-Take-away 10: CG persists across different MLLM backbones (RQ 5).
+(2025) have shown that jointly training classification and detection tasks can mutually enhance their performance.
+Building on this, we investigate whether MLLMs can leverage classification data (e.g., visual knowledge) and detection data (e.g., spatial information) through CG to improve downstream classification (Q1) or detection tasks (Q2).
+6.1 Experiment Settings
+Training Setup
+Each generalization combination used for training in this experiment includes one detection dataset and one classification dataset to examine the generalization relationship between these two vision tasks.
+The detailed training parameters can be found in Appendix A.6.
+Model Selection
+Next-Chat Zhang et al.
+(2023a) and MiniGPT-v2 Chen et al.
+(2023a) are selected as baselines, representing the two main approaches MLLMs use for detection tasks.
+The former treats bounding boxes as embeddings and decodes them into coordinates using a visual decoder, while the latter processes coordinate points as special text tokens and generates bounding box coordinates directly as output text.
+Data Processing
+Med-MAT includes both detection and segmentation datasets.
+If a segmentation dataset provides object localization using masks, we extract the outermost coordinates of the corresponding mask to construct a bounding box, facilitating generalization experiments for detection.
+Subsequently, to streamline the experiments, we structured the dataset following the official data formats of Next-Chat and MiniGPT-v2.
+6.2 Benefits for Classification (Q1)
+In this experiment, all possible CG combinations were selected and the CG-trained model will be tested on classification task.
+The final results in Figure 7 show that all CG combinations demonstrated the model’s successful utilization of detection data for CG to the Target data.
+6.3 Benefits for Detection (Q2)
+Subset 38 and 39 are selected as the objects in these datasets are relatively randomly distributed in the images, making them suitable for evaluating the model’s detection capability.
+Subsequently, we selected certain classification datasets to construct CG for testing and used cIoU to evaluate the detection performance (follow Chen et al.
+(2023a)).
+Since both baselines lack localization capabilities for medical tasks, we incorporated a fixed amount of Target data into our experiments, adjusting the evaluation scenario to assess support in low-data settings.
+The results in Table 5 show that all selected CG combinations help MLLMs achieve better performance in detection tasks.
+| Related Combination | Target Subset | Next-Chat | MiniGPT-v2 | |
+| - Skin | - Intestine | - Intestine | ||
+| - Intestine | - Skin | - Skin | ||
+Take-away 9: MLLMs can perform CG across classification and detection tasks.
 7 Related Work
-Medical MLLMs: Recently, adapting MLLMs to medical tasks has gained prominence due to their success in capturing complex visual features.
+Medical MLLMs Recently, adapting MLLMs to medical tasks has gained prominence due to their success in capturing complex visual features.
 Current MLLMs typically pair a visual encoder with a text-only LLM, aligning image data with language understanding.
@@
 (2024b) utilized specialized datasets and instruction tuning to refine medical VQA tasks.
-Generalization on Medical Imaging: Generalization in medical imaging Matta et al.
+Generalization on Medical Imaging
+Generalization in medical imaging Matta et al.
 (2024) has been extensively studied.
@@
 (2024) explore multiple aspects of medical image processing, examining how classification and segmentation tasks can mutually benefit each other.
-Detection with MLLMs: MiniGPT-4 Zhu et al.
-(2023) and LLaVA Liu et al.
-(2023) enhance LLMs by fine-tuning models on synthetic multimodal instruction data.
-However, these models are limited to image and text inputs and cannot generate detection bounding boxes or segmentation masks.
-Recent works have addressed these limitations using various strategies, such as encoding regions as features to allow models to accept regions as input Zhang et al.
+Detection with MLLMs
+Recent studies employ various strategies to equip MLLMs with the capability to handle detection tasks, such as encoding regions as features to allow models to accept regions as input Zhang et al.
 (2023b), representing object bounding box coordinates with text tokens Wang et al.
@@
 8 Conclusion
-To explore whether MLLMs can leverage CG to generalize unseen medical data, we constructed the Med-MAT dataset as the research platform for generalization experiments.
-The results confirmed the existence of CG and identified it as one of the primary factors behind the generalization capability of MLLMs in multi-task learning.
-Subsequent experiments demonstrated that increasing the volume of CG combinations consistently improved their effectiveness.
-Furthermore, CG enables medical tasks to efficiently fit with limited data, reducing dependence on the volume of training data.
-Importantly, CG is present across various MLLM backbones and can even facilitate generalization using data from tasks such as detection, underscoring its broad applicability.
+To investigate whether MLLMs can leverage CG to generalize to unseen medical data, we constructed the Med-MAT dataset as a research platform for generalization experiments.
+The results confirmed the presence of CG and identified it as a key factor of MLLMs’ generalization observed in multi-task learning.
+Further experiments showed that CG helps MLLMs handle limited data conditions, providing support for low-data medical tasks.
+Additionally, our findings showed that MLLMs can apply CG across detection and classification tasks, underscoring its broad generalization potential.
 Limitations
-The experiment confirms that MLLMs can use CG to understand unseen medical images and achieve data-efficient training.
-However, as shown in the results in 3.2, even after disrupting CG, the model’s generalization performance declines but still maintains some level of effectiveness.
-Thus, CG represents only one form of generalization for MLLMs in medical imaging.
-To support the exploration of other generalization mechanisms, we will make datasets publicly available according to the license.
-This study is focused exclusively on medical scenarios, but we believe that similar generalization effects may also apply in other multimodal tasks.
-Additionally, more granular medical generalization strategies, such as leveraging existing pneumonia data to generalize the detection of newly emerging types of pneumonia, are also promising areas for further investigation.
+The experiment confirms that MLLMs leverage CG for unseen medical images and data-efficient training.
+However, as shown in Section 4, disrupting CG reduces generalization but retains some effectiveness, indicating CG is just one aspect of MLLM generalization in medical imaging.
 Potential Risks
@@
 However, further experiments are needed to mitigate potential risks when deploying this concept in real-world medical settings.
+Acknowledgments
+This work was supported by Shenzhen Medical Research Fund (No.C2406002) from the Shenzhen Medical Academy of Research and Translation (SMART), the Shenzhen Science and Technology Program (JCYJ20220818103001002), Shenzhen Doctoral Startup Funding (RCBS20221008093330065), Tianyuan Fund for Mathematics of National Natural Science Foundation of China (NSFC) (12326608), Shenzhen Science and Technology Program (Shenzhen Key Laboratory Grant No.
+ZDSYS20230626091302006), and Shenzhen Stability Science Program 2023.
 Appendix A More Experiments
-A.1 CG on Detection Tasks (RQ 4)
-| Related Combination | Target Subset | Baseline | Trained | |
-| Lung, Lung Det | Bones, Diseases | Lung, Diseases | 49 | 52 |
-| Lung, Lung Det | Breast, Diseases | Lung, Diseases | 49 | 54 |
-| Bones, Spinal Error Det | Breast, Diseases | Bones, Diseases | 20 | 30 |
-| Bones, Spinal Error Det | Lung, Diseases | Bones, Diseases | 20 | 33 |
-| MRI, Diseases Det | End, Level | End, Diseases | 24 | 27 |
-| X-ray, Lung Det | CT, COVID | X-ray, COVID | 23 | 26 |
-| Der, Skin, Cancer Det | FP, Fundus, Diseases | Der, Skin, Diseases | 24 | 29 |
-| Mic, Cell, Cancer Det | CT, Kidney, Diseases | Mic, Cell, Diseases | 24 | 26 |
-| Related Combination | Target Subset | Baseline | Trained | |
-| Lung, Lung Det | Bones, Diseases | Lung, Diseases | 41 | 47 |
-| Lung, Lung Det | Breast, Diseases | Lung, Diseases | 41 | 49 |
-| Bones, Spinal Error Det | Breast, Diseases | Bones, Diseases | 31 | 35 |
-| Bones, Spinal Error Det | Lung, Diseases | Bones, Diseases | 31 | 37 |
-| MRI, Diseases Det | End, Level | End, Diseases | 24 | 26 |
-| X-ray, Lung Det | CT, COVID | X-ray, COVID | 22 | 23 |
-| Der, Skin, Cancer Det | FP, Fundus, Diseases | Der, Skin, Diseases | 27 | 30 |
-| Mic, Cell, Cancer Det | CT, Kidney, Diseases | Mic, Cell, Diseases | 20 | 24 |
-Previous studies have shown that training on detection datasets can improve model performance in classification tasks.
-Building on this insight, we further examine its applicability to CG.
-For this investigation, all possible detection subset combinations in Med-MAT were selected to explore their impact on CG (Selection Strategy).
-Experimental Setup: We conducted generalization experiments for detection and classification.
-Specifically, we performed generalization validation on Next-Chat Zhang et al.
-(2023a) and MiniGPT-v2 Chen et al.
-(2023a).
-Next-Chat models the bounding box as an embedding and utilizes a decoder for decoding, while MiniGPT-v2 treats the bounding box as a text token, which are common approaches used by existing MLLM implementations for detection.
-By conducting CG validation using distinct bounding box modeling methods, we further demonstrate the broad applicability of the CG approach.
-Each experiment was conducted on 8 A800 (80GB) GPUs.
-The two backbone was trained separately in this experiment.
-For Next-Chat, we directly trained the model in its second training stage and fine-tuned it for 2 epochs on our composition datasets with a learning rate of 2e-5, keeping all other training parameters at their default settings.
-Similarly, for MiniGPT-v2, we trained the backbone model from the second stage, starting with a learning rate of 2e-5 and gradually reducing it to 2e-6 over 3 epochs.
-The experiments in Table 5 and Table 6 present the results for Next-Chat and MiniGPT-v2, respectively.
-The separation regions closely align with the 4 Direction Types defined in the classification task, except for Fix Task, which is excluded due to the substantial differences between detection and classification tasks, making it impossible to find the fix one.
-The results demonstrate that joint training on detection and classification tasks effectively compositional generalizes to classification tasks in both backbones, leading to higher model accuracy.
-This indicates that CG is also present in these tasks.
-| Related Combination | Target Subset | Baseline | Trained | |
-| Bones, State | Breast, Diseases | Bones, Diseases | 61 | 65 |
-| Lung, COVID | Bones, Diseases | Lung, Diseases | 80 | 91 |
-| CT, COVID | X-ray, Diseases | X-ray, COVID | 35 | 40 |
-| CT, State | X-ray, Diseases | X-ray, State | 35 | 43 |
-| X-ray, Lung | CT, Brain(Cancer) | CT, Lung | 32 | 33 |
-| X-ray, Lung | CT, Brain | CT, Lung(Cancer) | 65 | 72 |
-| FP, Fundus, Diseases | Mic, Cell, Level | FP, Fundus, Level | 48 | 45 |
-| Mic, Cell, Cell Identification | FP, Fundus, Level | Mic, Cell, Level | 34 | 41 |
-| Related Combination | Target Subset | Baseline | Trained | |
-| Bones, State | Breast, Diseases | Bones, Diseases | 52 | 59 |
-| Lung, COVID | Bones, Diseases | Lung, Diseases | 64 | 75 |
-| CT, COVID | X-ray, Diseases | X-ray, COVID | 33 | 38 |
-| CT, State | X-ray, Diseases | X-ray, State | 33 | 41 |
-| X-ray, Lung | CT, Brain(Cancer) | CT, Lung | 31 | 29 |
-| X-ray, Lung | CT, Brain | CT, Lung(Cancer) | 49 | 57 |
-| FP, Fundus, Diseases | Mic, Cell, Level | FP, Fundus, Level | 55 | 61 |
-| Mic, Cell, Cell Identification | FP, Fundus, Level | Mic, Cell, Level | 10 | 32 |
-A.2 Other MLLMs (RQ 5)
-The choice to use models like LLaVA, MiniGPT-v2 and NExT-GPT is because their training data and overall process are publicly available, ensuring that they have not been too much exposed to medical images, thus preventing bias in the medical image knowledge injection into the MLLM.
+A.1 Benefits for Segmentation
+Segmentation-enabled LLMs, such as Next-GPT, first use the LLM to identify potential regions of the target object and then apply a SAM to decode the object mask, thereby completing the segmentation task.
+In this context, segmentation can be seen as an extension of detection, potentially requiring more images to achieve improved performance.
+We conducted additional experiments to explore whether MLLMs can still utilize CG to understand new images across both segmentation and classification tasks.
+| Related Combination | Target Subset | Next-Chat | |
+| - Skin | - Intestine | - Intestine | |
+| - Intestine | - Skin | - Skin | |
+The results in Table 6 demonstrate that, in the context of segmentation tasks, MLLMs are still able to leverage CG to understand new tasks, which is consistent with our original conclusions.
+A.2 More Complex Medical Elements
+While MAT-Triplet Categorization is useful, predefined categories may limit the exploration of more complex medical attributes, so we also considered integrating more flexible categorization to explore additional medical attributes.
+Additional Element 1: Population Groups
+We selected VinDr-PCXR and MedMAT Subset 31 for the experiment, as they contain X-ray images of children and adult groups, respectively. The results are shown in Table 7.
+Additional Element 2: Finer Disease
+"Finer disease" means more detailed categorization.
+For instance, we treat COVID and common pneumonia as distinct diseases for generalization.
+We split the Normal data in the training set into two parts and combined each with COVID and Pneumonia data to create new datasets.
+The results are shown in Table 8.
+| Related Combination | Target Subset | LLaVA | ||||
+Children (CG) |
+| Related Combination | Target Subset | LLaVA | ||||
+| Related Combination | Target Subset | Baseline | 1st | 2nd | 3rd | Mean and SD | ||||
+A.3 Statistical Tests of the Generalization Results
+To ensure consistency and repeatability of the experiment, we performed statistical tests in this section.
+LLaVA is selected as the baseline, and we used the same data combinations from Section 3.3.
+Each experiment was repeated 3 times, and we reported the mean and standard deviation (SD) of the results.
+From the results in Table 9, we can observe that the outcomes across runs show low variance, indicating overall stability, and they continue to support our original experimental conclusions.
+A.4 CG with All MAT-Triplet Elements from Different Sources
+In previous controlled experiments (Section 3), one element of the MAT-Triplet was kept constant while CG was explored in the remaining two elements.
+To ensure that all the 3 MAT-Triplet elements of the target data originated from three distinct datasets, additional experiments were conducted to further validate the effectiveness of CG.
+For these experiments, all possible combinations meeting the criteria in Med-MAT were selected (Selection Strategy).
+The results presented in Table 10 demonstrate that most combinations can effectively generalize to the Target data.
+| Related Combination | Target Subset | Baseline | Trained | CG Helps | ||||
+28 | 26 | ✗ |
+28 | 25 | ✗ |
+33 | 64 | ✓ |
+33 | 70 | ✓ |
+30 | 45 | ✓ |
+30 | 38 | ✓ |
+30 | 44 | ✓ |
+31 | 32 | ✓ |
+31 | 52 | ✓ |
+| Related Combination | Target Subset | Baseline | Trained | CG Helps | ||||
+61 | 65 | ✓ |
+80 | 91 | ✓ |
+35 | 40 | ✓ |
+35 | 43 | ✓ |
+32 | 33 | ✓ |
+65 | 72 | ✓ |
+48 | 45 | ✗ |
+34 | 41 | ✓ |
+| Related Combination | Target Subset | Baseline | Trained | CG Helps | ||||
+52 | 59 | ✓ |
+64 | 75 | ✓ |
+33 | 38 | ✓ |
+33 | 41 | ✓ |
+31 | 29 | ✗ |
+49 | 57 | ✓ |
+55 | 61 | ✓ |
+10 | 32 | ✓ |
+Analysis of the results
+The results in Table 7 and 8 indicate that the two new attributes show data leakage due to subtle visual differences in corresponding images (e.g., COVID-19 and pneumonia have similar features).
+Importantly, the MLLM trained with CG combinations still shows improvements on downstream tasks, confirming that our approach remains valid for new attributes.
+Reason to choose the existing three attributes (MAT-Triplet: Modality, Area, Task)
+We have considered additional categories such as age, gender, and finer disease classification, but we ultimately chose to focus on the MAT-Triplet categories for the following reasons.
+The boundaries between MAT-Triplet (Modality, Area, Task) are clear.
+Different modalities and areas correspond to distinct imaging methods and body areas, leading to significant differences between images; different tasks also require the MLLM to extract specific information, demanding varied understanding of the images.
+All datasets can be annotated using MAT-Triplet (Modality, Area, Task) easily.
+Other medical labels, such as gender and age, are only available in a small portion of datasets and are not suitable for large-scale annotation.
+Similar categorization strategies have been adopted in previous studies.
+| Related Combination | Target Subset | Baseline | Trained | CG Helps | ||||
+49 | 52 | ✓ |
+49 | 54 | ✓ |
+20 | 30 | ✓ |
+20 | 33 | ✓ |
+24 | 27 | ✓ |
+23 | 26 | ✓ |
+24 | 29 | ✓ |
+24 | 26 | ✓ |
+| Related Combination | Target Subset | Baseline | Trained | CG Helps | ||||
+41 | 47 | ✓ |
+41 | 49 | ✓ |
+31 | 35 | ✓ |
+31 | 37 | ✓ |
+24 | 26 | ✓ |
+22 | 23 | ✓ |
+27 | 30 | ✓ |
+20 | 24 | ✓ |
+A.5 Details of Section 3.3: Exploring CG on different MLLM Backbones
 To ensure the experiment results are not influenced by the model choice, we also tested several other models on some subsets of Med-MAT and observed similar results.
@@
 The training was performed with a batch size of 32 and a learning rate set to 2e-6, ensuring that all parameters were updated to optimize the model performance.
-The experimental results presented in Table 7 and Table 8 demonstrate that the CG persists across these different backbone architectures.
-This observation indicates that CG is not confined to a specific model type, thereby highlighting its universal applicability and robustness across various model frameworks.
-Such findings underscore the versatility of CG, suggesting it can be effectively integrated into a wide range of models.
+| Related Combination | Target Subset | HuatuoGPT | ||||
+A.6 Details of Section 6: Exploring CG across Detection and Classification
+Experimental Setup: We conducted generalization experiments for detection and classification.
+Specifically, we performed generalization validation on Next-Chat Zhang et al.
+(2023a) and MiniGPT-v2 Chen et al.
+(2023a).
+Next-Chat models the bounding box as an embedding and utilizes a decoder for decoding, while MiniGPT-v2 treats the bounding box as a text token, which are common approaches used by existing MLLM implementations for detection.
+By conducting CG validation using distinct bounding box modeling methods, we further demonstrate the broad applicability of the CG approach.
+Each experiment was conducted on 8 A800 (80GB) GPUs.
+The two backbones were trained separately in this experiment.
+For Next-Chat, we directly trained the model in its second training stage and fine-tuned it for 2 epochs with a learning rate of 2e-5, keeping all other training parameters at their default settings.
+Similarly, for MiniGPT-v2, we trained the backbone model from the second stage, starting with a learning rate of 2e-5 and gradually reducing it to 2e-6 over 3 epochs.
+A.7 CG with Medical Multimodal LLM
+In previous experiments, general MLLMs are selected to prevent the MLLM’s inherent medical knowledge from affecting CG results.
+Our experiments focus on how MLLMs leverage CG to interpret unseen medical images.
+If the model has learned some fundamental elements of the Target data, it would compromise the fairness of the experiments.
+To demonstrate that our results still work on medical LLMs, we employed the same data combinations from Section 3.3 to investigate CG on medical MLLMs (we selected HuatuoGPT-Vision as the baseline).
+The results in Table 15 demonstrate that the medical-expert MLLM can still leverage CG to enhance their performance on novel tasks, further supporting the validity and consistency of our findings.
 Appendix B The Dataset: Med-MAT
@@
 We hope this dataset will support and advance future generalization experiments on medical imaging.
+B.4 Data Sources and Distribution
+All Med-MAT data are sourced from public medical image challenges or widely used, high-impact datasets previously applied in deep learning training, ensuring reliable annotations.
+Before inclusion in Med-MAT, all datasets underwent label averaging where possible; test sets, in particular, were strictly balanced to ensure accuracy reliably reflects model performance.
+Each Med-MAT training subset contains 3,000 samples, while test sets maximize size under label balance constraints.
+Appendix C Bad cases analysis and solutions
+C.1 Bad case analysis
+Some Trained models show minimal gains or even performance declines in Table 3, with classification accuracy lower than either the Baseline or Baseline+.
+After a thorough examination, we found that these Target datasets require more fine-grained medical condition classification.
+Beyond disease presence, they need detailed assessments, such as severity grading (e.g., bone age estimation, cancer staging) or distinguishing similar conditions (e.g., differentiating COVID-19 from pneumonia).
+The Related combinations lack suitable fundamental elements: For CG, the training data must include the Target task’s core elements.
+Here, we use other "level classification/grading" tasks for generalization, but their criteria differ significantly, misaligning with the Target task’s needs.
+Without defined grading standards, MLLMs lacking relevant knowledge can’t perform fine-grained tasks: Tasks like bone age assessment and cancer staging vary by criteria, and without this knowledge, MLLMs can’t accurately classify them.
+C.2 Possible solutions
+Few-shot prompting
+As we illustrated before, most of the bad cases involve fine-grained tasks needing specialized knowledge.
+So, in order to minimize the effect of a lack of relevant knowledge, we also conducted few-shot experiments to add some target images in the prompts.
+Subset X-ray, Lung, Normal-COVID-Pneumonia was chosen for its simple structure, with LLaVA as the baseline.
+We randomly sampled n images per label for n-shot inference and repeated each experiment 3 times.
+| Model | 0-shot | 2-shot | 3-shot | 4-shot |
+| LLaVA | ||||
+| LLaVA + CG |
+The results in Table 16 demonstrate that training with CG combinations can improve the few-shot performance of MLLMs on downstream tasks, even when direct CG generalization is not effective.
+Adding some Target data in training
+As described in Section 5.2, we selected cases where CG alone couldn’t achieve satisfactory results and augmented their training sets with target data.
+The results in this section indicate that while CG may not directly enhance generalization, it accelerates the model’s adaptation to downstream tasks.
 | Subset No. | Modality | Anatomical Area | Task | Datasets No. |
@@
 | 3 | SARS-COV-2 Ct-Scan | COVID19, Classification Dataset | Soares et al. (2020) |
-| 4 | COVID CT COVID-CT | COVID19, Classification Dataset. | Zhao et al. (2020) |
+| 4 | COVID CT COVID-CT | COVID19, Classification Dataset | Zhao et al. (2020) |
 | 5 | Chest CT-Scan | Cancer Classification | SunneYi (2021) |
@@
 | 10 | ISIC 2020 | Melanoma, Benign or Malignant | Rotemberg et al. (2021) |
-| 11 | PED-UFES-20 | Skin Multi Classification | Pacheco et al. (2020) |
+| 11 | PAD-UFES-20 | Skin Multi Classification | Pacheco et al. (2020) |
 | 12 | Web-scraped Skin Image | Skin Desease Multi Classification | Islam et al. (2022b) |
@@
 | 28 | Adam dataset | Age-related Macular Degeneration | Liang (2021) |
-| 29 | APTOS 2019 Blindness | Blindness Level Identification 0 4 | Karthik et al. (2019) |
-| 30 | DRIMBD | Quality Testing of Retinal Images | Prentasic et al. (2013) |
+| 29 | APTOS 2019 Blindness | Blindness Level Identification | Karthik et al. (2019) |
+| 30 | DRIMDB | Quality Testing of Retinal Images | Prentasic et al. (2013) |
 | 31 | Glaucoma Detection | Glaucoma Classification | Zhang and Das (2022) |
@@
 | 34 | SICAPv2 | Cancer Degree Classification | Silva-Rodríguez et al. (2020) |
-| 35 | Blood Cell Images | Blood Cell Classificaion (Multi) | Mooney (2017) |
+| 35 | Blood Cell Images | Blood Cell Classificaion | Mooney (2017) |
 | 36 | BreakHis | Cell type and beginormag | Bukun (2019) |
-| 37 | Chaoyang | Multi Classification of pathologists | Zhu et al. |
+| 37 | Chaoyang | Multi Classification of pathologists | Zhu et al. (2021a) |
 | 38 | HuSHeM | Sperm Head Morphology Classificaion | Shaker (2018) |
@@
 | 53 | Knee Osteoarthritis Dataset | Knee Osteoarthritis with severity grading | Chen (2018) |
-| 54 | Shenzhen Chest X-Ray Set | COVID19, Classification Dataset. | Jaeger et al. (2014) |
+| 54 | Shenzhen Chest X-Ray Set | COVID19, Classification Dataset | Jaeger et al. (2014) |
 | 55 | Chest X-ray PD | COVID and Pneumonia | Asraf and Islam (2021) |
@@
 | 69 | Arthrosis | Bone Age Classification | Zha (2021) |
-| 70 | HSA-NRL | Multi Classification of pathologists | Zhu et al. (2021) |
+| 70 | HSA-NRL | Multi Classification of pathologists | Zhu et al. (2021b) |
 | 71 | ISIC 2018 (Task 3) | Multi Classification | Codella et al. (2019) |
@@
 | 83 | PBC dataset normal DIB | Multi Classification | Acevedo et al. (2020) |
-| 84 | Human Protein Atlas | Multi-label Classification (Only green) | Le et al. (2022) |
+| 84 | Human Protein Atlas | Multi-label Classification | Le et al. (2022) |
 | 85 | RSNA Pneumonia Detection Challenge 2018 | Multi-label Classification | Anouk Stein et al. (2018) |